### v2

# Feynman Graph Integrals on Kähler Manifolds

Abstract
In this paper, we provide a rigorous proof of the convergence, in the Cauchy principal value sense, of Feynman graph integrals of holomorphic quantum field theories on closed real analytic Kähler manifolds. In the special case of Calabi–Yau manifolds, these graph integrals give rise to new invariants of Calabi–Yau metrics.
1 Introduction
Feynman graph integrals, originally introduced in physics, have become central to numerous areas of mathematics. A notable example are the configuration space integrals for topological quantum field theories developed by S. Axelrod and I. Singer [2, 3] and by M. Kontsevich [21]. These integrals have proved exceptionally powerful, forming the foundation of modern approaches to knot invariants [21], 3d manifold invariants [7, 6], smooth structures on fiber bundles [21], operad theory [16, 23], deformation quantization of Poisson manifolds [24], and many other applications.
In contrast, a systematic theory of Feynman graph integrals in holomorphic quantum field theories has only begun to take shape in the past decade [12, 25, 27, 28, 26, 8, 36, 34, 33]. A key technical obstacle is that these integrals are typically not absolutely convergent (see [33, Example 2.1.1]). To address this, the first author recently introduced a heat-kernel regularization and established the finiteness of holomorphic field theories on affine spaces; see [33]. See also [11] for further discussion of heat-kernel regularization in renormalization in quantum field theory.
This paper presents an alternative approach to [33], offering another formulation of Feynman graph integrals in holomorphic quantum field theories using Cauchy principal values (see [18]). Our construction extends the setting of [33] from affine spaces to arbitrary closed real analytic Kähler manifolds, thereby broadening the scope of holomorphic quantum field theories.
1.1 Main results
For simplicity, and to avoid introducing too many concepts and notations at the beginning of the paper, we state here a special case of our main results (Theorem 2.20 and Theorem 2.25).
Let $M$ be a closed Calabi-Yau manifold, whose holomorphic volume form is denoted by $\Omega$.
Let $P$ be a distribution-valued section of the bundle $\Lambda^{\bullet}T_{0,1}^{*}(M\times M)$ over $M\times M$, such that the following equations hold (see Definition 2.5 for details):
$\begin{cases}\left(\bar{\partial}\otimes\operatorname{id}+\operatorname{id}\otimes\bar{\partial}\right)P=\delta-\mathcal{H},\\ P\in\mathrm{Im}(\bar{\partial}^{*}\otimes\mathrm{id}),\end{cases}$
where $\delta$ and $\mathcal{H}$ are the integral kernels of the identity operator and the harmonic projection operator, respectively.
Let $\tilde{\rho}:M\times M\rightarrow\mathbb{R}_{\geq 0}$ be a non-negative function, such that:
  1. (1)
    $\tilde{\rho}^{2}$ is smooth and $\tilde{\rho}^{-1}(0)=\triangle$, where $\triangle:=\{(p,p)\in M\times M:p\in M\}$.
  2. (2)
    There exists an open neighborhood $U\subset M\times M$ of $\triangle$ such that
    $\tilde{\rho}^{2}|_{U}=\rho^{2}|_{U},$
    where $\rho$ is the distance function on $M$.
Then we have
Theorem 1.1.
Let $\vec{\Gamma}$ be a directed graph, whose vertex set $\vec{\Gamma}_{0}$ and edge set $\vec{\Gamma}_{1}$ are ordered sets. Then
$\lim_{\epsilon\rightarrow 0}\int_{M^{\vec{\Gamma}_{0}}}\chi_{\epsilon}(p_{1},\cdots,p_{|\vec{\Gamma}_{0}|})\Big{(}\prod_{e\in\vec{\Gamma}_{1}}P\big{(}p_{h(e)},p_{t(e)}\big{)}\Big{)}\Big{(}\prod_{i\in\vec{\Gamma}_{0}}\Omega(p_{i})\Big{)}$
exists, where $t,h:\vec{\Gamma}_{1}\rightarrow\vec{\Gamma}_{0}$ denotes the tail and head of a given directed edge, and $|\vec{\Gamma}_{0}|$ is the number of vertices. Here $\chi_{\epsilon}$ is the function on $M^{|\vec{\Gamma}_{0}|}$ defined by
$\chi_{\epsilon}(p_{1},\cdots,p_{|\vec{\Gamma}_{0}|})=\begin{cases}0,&\text{if }\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}(p_{i},p_{j})\leq\epsilon,\\ 1,&\text{if }\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}(p_{i},p_{j})>\epsilon.\end{cases}$
Moreover, the integral does not depend on the choice of $\tilde{\rho}$.
By the uniqueness of $P$ (see 2.6), it follows that the integrals defined in 1.1 are invariants of the Calabi–Yau metric and the holomorphic volume form.
Using Cauchy principal values in Appendix B and the Fulton–MacPherson compactification of the configuration spaces, we can reformulate Theorem 1.1 as follows:
Theorem 1.2.
Let $\vec{\Gamma}$ be a directed graph, whose vertex set $\vec{\Gamma}_{0}$ and edge set $\vec{\Gamma}_{1}$ are ordered sets. Then
$\lim_{\epsilon\rightarrow 0}\int_{M^{\vec{\Gamma}_{0}}}\chi_{\epsilon}(p_{1},\dots,p_{|\vec{\Gamma}_{0}|})\Big{(}\prod_{e\in\vec{\Gamma}_{1}}P(p_{h(e)},p_{t(e)})\Big{)}\Big{(}\prod_{i\in\vec{\Gamma}_{0}}\Omega(p_{i})\Big{)}$
$=\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{\widetilde{Conf}_{\vec{\Gamma}_{0}}(M)}\Big{(}\prod_{e\in\vec{\Gamma}_{1}}P(p_{h(e)},p_{t(e)})\Big{)}\Big{(}\prod_{i\in\vec{\Gamma}_{0}}\Omega(p_{i})\Big{)},$
where $\widetilde{Conf}_{\vec{\Gamma}_{0}}(M)$ is the Fulton–MacPherson compactification of the configuration space of $|\vec{\Gamma}_{0}|$ points in $M$ (see 2.14 ), and the integral on the right hand side is in the sense of Cauchy principal values (see Appendix B).
Theorem 1.1 and Theorem 1.2 are special cases of our main results Theorem 2.20 and Theorem 2.25, where we consider general closed real analytic Kähler manifolds and Hermitian holomorphic vector bundles with holomorphic pairings.
Notably, the proof of our main theorems draws on Getzler’s rescaling technique (see [4, 15]), originally developed for the Atiyah–Singer local index theorem, highlighting a fruitful interplay between index-theoretic methods and Feynman graph integrals. In future work, we plan to adapt the machinery employed in the proof of the family version of the Atiyah–Singer local index theorem to the study of Feynman graph integrals.
1.2 Applications and outlook
Similar to topological quantum field theories, the Feynman graph integrals of holomorphic field theories have numerous applications. Below, we highlight a few:
  1. (1)
    Constructions of geometric invariants as Theorem 1.1.
  2. (2)
    Construction of factorization algebras, as in [34]; see also [9, 10] for foundational discussions on factorization algebras.
  3. (3)
    Constructions of higher chiral algebras as in [17].
Moreover, the results established in this paper have potential applications in the following areas:
  1. (1)
    Gauge field theories on $\mathbb{R}^{4}$. In the well-known Penrose–Ward correspondence (see [31, 35]), it is conjectured that perturbative self-dual Yang-Mills theory should correspond to a perturbative holomorphic quantum field theory on the twistor space. We expect that our theorems offer new insights toward understanding this correspondence. For example, our results should imply the finiteness of Feynman graph integrals in self-dual Yang-Mills theory.
  2. (2)
    Holomorphic “knot” invariants. Feynman graph integrals arising from topological field theories have proven useful in the study of knot invariants. Our theorems have potential applications in the study of embeddings of Riemann surfaces—viewed as holomorphic “knots”—into complex manifolds. For the study of holomorphic linking numbers in the literature, see [1, 13].
  3. (3)
    Mirror Symmetry. In mirror symmetry, the Gromov-Witten invariants of Calabi-Yau manifolds are conjecturally equal to the Feynman graph integrals of Kodaira-Spencer gravity theory (or BCOV theory, see [5]) on the mirror manifolds. Since Kodaira-Spencer gravity theory is a holomorphic quantum field theory, it is natural to expect that our theorems can be applied in this context.
1.3 Organizations
In Section 2, we introduce the notion of the propagator. Assuming 2.11—that is, that the propagator has the desired singularities, namely (logarithmic) divisorial type singularities (B.12)—we define Feynman graph integrals using Cauchy principal values. This construction forms the core of our main results, Theorem 2.20 and Theorem 2.25.
Section 3 and Section 4 are devoted to the proof of 2.11. In Section 3, we introduce the notion of regular expressions, which can be viewed as the analog of divisorial type singularities in the context of heat kernels. In Section 3.2, we establish the connection between regular expressions of the heat kernel and the divisorial type singularities of the propagator, as stated in 3.14. This reduces the proof of 2.11 to verifying that the heat kernel admits a regular expression, namely 4.3.
Finally, in Section 4, we prove 4.3 using Getzler’s rescaling technique.
Acknowledgments
We thank Keyou Zeng, Zhengping Gui, Si Li, Ezra Getzler, Maciej Szczesny, Brian Williams, and Kevin Costello for valuable discussions. Part of this work was completed during the first author’s visit to the Institut Mittag‐Leffler. He gratefully acknowledges its hospitality.
2 Feynman Graph Integrals on Kähler Manifolds
In this section, we introduce the Feynman graph integrals from holomorphic quantum field theories.
2.1 Propagators
In this subsection, we introduce the notion of the propagator in holomorphic quantum field theory, and establish its existence (2.8) and uniqueness (2.6).
We begin by recalling some basic facts from Hodge theory on Kähler manifolds.
Let $M$ be a closed Kähler manifold, and let $E\to M$ be a Hermitian holomorphic vector bundle.
[…]
$\Delta_{\widetilde{E}}=\bar{\partial}_{\widetilde{E}}\circ\bar{\partial}_{\widetilde{E}}^{*}+\bar{\partial}_{\widetilde{E}}^{*}\circ\bar{\partial}_{\widetilde{E}}$.
[…]
$(s_{1}\otimes u_{1},\,s_{2}\otimes u_{2})=\int_{M}\langle s_{1},s_{2}\rangle\,u_{1}\wedge u_{2},$ (2.1)
where $s_{1}\in\Gamma(E^{*})$, $s_{2}\in\Gamma(E)$, $u_{1}\in\Gamma(K_{M}\otimes\Lambda^{\bullet}T_{0,1}^{*}M)$, and $u_{2}\in\Gamma(\Lambda^{\bullet}T_{0,1}^{*}M)$, and ${\langle}s_{1},s_{2}\rangle$ denotes the natural pairing between $E$ and $E^{*}$.
Remark 2.1.
In this paper, $\Gamma(E)$ denotes the space of smooth sections of a vector bundle $E$, rather than holomorphic sections. When necessary, we also write $\Gamma(E)$ as $\Gamma(M;E)$ to emphasize the base manifold $M$.
Lemma 2.2.
[…]
  1. (1)
    $(\bar{\partial}_{\widetilde{E}^{\prime}}\alpha,\beta)=(-1)^{|\alpha|+1}(\alpha,\bar{\partial}_{\widetilde{E}}\beta).$
  2. (2)
    $(\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha,\beta)=(-1)^{|\alpha|}(\alpha,\bar{\partial}_{\widetilde{E}}^{*}\beta).$
Proof.
The first assertion follows from Stokes’ theorem.
To prove the second, recall the identity:
$\bar{\partial}_{\widetilde{E}}^{*}=-\bar{*}_{E^{*}}\circ\bar{\partial}_{\widetilde{E}^{\prime}}\circ\bar{*}_{E},$ (2.2)
[…]
$\quad(\alpha,\bar{\partial}_{\widetilde{E}}^{*}\beta)=-(\alpha,\bar{*}_{E^{*}}\bar{\partial}_{\widetilde{E}^{\prime}}\bar{*}_{E}\beta)=-(\alpha,\bar{\partial}_{\widetilde{E}^{\prime}}\bar{*}_{E}\beta)_{\widetilde{E}^{\prime}}$
$=-\overline{(\bar{\partial}_{\widetilde{E}^{\prime}}\bar{*}_{E}\beta,\alpha)_{\widetilde{E}^{\prime}}}=-\overline{(\bar{\partial}_{\widetilde{E}^{\prime}}\bar{*}_{E}\beta,\bar{*}_{E^{*}}\alpha)}$
$=(-1)^{|\beta|}\overline{(\bar{*}_{E}\beta,\bar{\partial}_{\widetilde{E}}\bar{*}_{E^{*}}\alpha)}=(-1)^{|\beta|+|\alpha|+1}\overline{(\bar{*}_{E}\beta,\bar{*}_{E^{*}}\bar{*}_{E}\bar{\partial}_{\widetilde{E}}\bar{*}_{E^{*}}\alpha)}$
$=(-1)^{|\beta|+|\alpha|}\overline{(\bar{*}_{E}\beta,\bar{*}_{E^{*}}\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha)}=(-1)^{|\beta|+|\alpha|}(\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha,\bar{*}_{E}\beta)_{\widetilde{E}}$
$=(-1)^{|\beta|+|\alpha|}(\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha,\bar{*}_{E^{*}}\bar{*}_{E}\beta)=(-1)^{|\alpha|}(\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha,\beta).$
This completes the proof. ∎
[…]
$\begin{cases}(\alpha,\bar{\partial}_{\widetilde{E}}\beta):=(-1)^{|\alpha|+1}(\bar{\partial}_{\widetilde{E}^{\prime}}\alpha,\beta),\\ (\alpha,\bar{\partial}_{\widetilde{E}}^{*}\beta):=(-1)^{|\alpha|}(\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha,\beta).\end{cases}$
By Hodge theory, we have the decomposition
[…]
Let $\mathcal{H}^{\prime}$ be the harmonic projection operator, defined as the composition
[…]
Consider the distribution-valued section $\mathcal{H}^{\prime\prime}$ of $\widetilde{E}\boxtimes\widetilde{E}^{\prime}$ given by
$(\alpha\boxtimes\beta,\mathcal{H}^{\prime\prime})=(\alpha,\mathcal{H}^{\prime}\beta),$
[…]
Lemma 2.3.
The distribution-valued section $\mathcal{H}^{\prime\prime}$ turns out to be a smooth section of $\widetilde{E}\boxtimes\widetilde{E}^{\prime}$.
Proof.
We note $\mathcal{H}^{\prime\prime}$ satisfies the following equations
$\begin{cases}(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}^{\prime}})\mathcal{H}^{\prime\prime}=0\\ (\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}^{\prime}}^{*})\mathcal{H}^{\prime\prime}=0.\end{cases}$
Therefore, $\mathcal{H}^{\prime\prime}$ is a harmonic section of $\widetilde{E}\boxtimes\widetilde{E}^{\prime}$. Since the Laplacian operator is an elliptic operator, we conclude that $\mathcal{H}^{\prime\prime}$ is a smooth section of $\widetilde{E}\boxtimes\widetilde{E}^{\prime}$ by the regularity of elliptic differential equations (see [19]). ∎
Now, we assume that there is a non-degenerate holomorphic bundle map
$\omega:E\otimes E\to K_{M}.$
The non-degeneracy induces two isomorphisms:
$\begin{cases}\omega_{L}:\alpha\in\Gamma(E)\rightarrow\omega(\alpha\otimes-)\in\Gamma(E^{*}\otimes K_{M})\\ \omega_{R}:\alpha\in\Gamma(E)\rightarrow\omega(-\otimes\alpha)\in\Gamma(E^{*}\otimes K_{M}).\end{cases}$
We denote the inverses of the induced isomorphisms by $\omega^{-1}_{L}$ and $\omega^{-1}_{R}$.
Definition 2.4.
[…]
$\mathcal{H}:=(\mathrm{id}\otimes\omega^{-1}_{L})\mathcal{H}^{\prime\prime}.$
[…]
$\delta(\alpha\boxtimes\beta)=\big{(}\alpha,\omega^{-1}_{R}(\beta)\big{)},$
[…]
Now, we can define propagators.
Definition 2.5.
Let $(M,E,\omega)$ be a triple as in 2.4. A propagator is a distributional section of $\widetilde{E}\boxtimes\widetilde{E}$, such that
[…]
The propagator exists and is unique. Let’s first prove the uniqueness.
Proposition 2.6.
[…]
Proof.
[…]
$(P-P^{\prime},\,\alpha\boxtimes\beta)=0.$
We begin by observing that
$\left\{\begin{aligned} &\big{(}\bar{\partial}_{\widetilde{E}}\otimes\operatorname{id}+\operatorname{id}\otimes\bar{\partial}_{\widetilde{E}}\big{)}\big{(}P-P^{\prime}\big{)}=0,\\ &P-P^{\prime}\in\operatorname{Im}(\bar{\partial}_{\widetilde{E}}^{*}\otimes\operatorname{id}).\end{aligned}\right$. (2.3)
[…]
$P-P^{\prime}=(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id})G.$
[…]
$\alpha=\alpha_{H}+(\bar{\partial}_{\widetilde{E}^{\prime}}\bar{\partial}_{\widetilde{E}^{\prime}}^{*}+\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\bar{\partial}_{\widetilde{E}^{\prime}})\alpha^{\prime},$
where $\alpha_{H}$ is harmonic. Then we compute:
$\quad(P-P^{\prime},\,\alpha\boxtimes\beta)=\big{(}P-P^{\prime},\,(\alpha_{H}+\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\bar{\partial}_{\widetilde{E}^{\prime}}\alpha^{\prime})\boxtimes\beta\big{)}+\big{(}P-P^{\prime},\,(\bar{\partial}_{\widetilde{E}^{\prime}}\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha^{\prime})\boxtimes\beta\big{)}$
$=\big{(}(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id})G,\,(\alpha_{H}+\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\bar{\partial}_{\widetilde{E}^{\prime}}\alpha^{\prime})\boxtimes\beta\big{)}\pm\big{(}(\bar{\partial}_{\widetilde{E}^{\prime}}\otimes\mathrm{id})(P-P^{\prime}),\,(\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha^{\prime})\boxtimes\beta\big{)}$
$=\pm\Big{(}G,\,\big{(}\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\otimes\mathrm{id}\big{)}\big{(}(\alpha_{H}+\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\bar{\partial}_{\widetilde{E}^{\prime}}\alpha^{\prime})\boxtimes\beta\big{)}\Big{)}\pm\big{(}(\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}^{\prime}})(P-P^{\prime}),\,(\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha^{\prime})\boxtimes\beta\big{)}$
$=0\pm\big{(}(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id})(\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}^{\prime}})G,\,(\bar{\partial}_{\widetilde{E}^{\prime}}^{*}\alpha^{\prime})\boxtimes\beta\big{)}=0.$
This completes the proof.
Now, we construct the propagator using the heat kernel.
Lemma 2.7.
Let $(M,E,\omega)$ be a triple as in 2.4. Then there exists a unique solution $)H_{t}\in C^{\infty}((0,\infty);\Gamma(\widetilde{E}\boxtimes\widetilde{E}$)) satisfying:
$\begin{cases}\displaystyle\partial_{t}H_{t}=-(\Delta_{\widetilde{E}}\otimes\mathrm{id})H_{t},\\ \displaystyle\lim_{t\to 0}H_{t}=\delta.\end{cases}$ (2.4)
Moreover, we have $\lim_{t\to\infty}H_{t}=\mathcal{H}$ and
$(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})H_{t}=0.$ (2.5)
Here $\delta$ and $\mathcal{H}$ are induced by the pairing $\omega$ as in 2.4.
Proof.
The existence and uniqueness of the heat kernel follow from [4]. The convergence $\lim_{t\to\infty}H_{t}=\mathcal{H}$ could be proved as in [4, Section 2.6].
To show that
$(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})H_{t}=0,$
we note that
$(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})H_{t}$
satisfies the heat equation:
$\begin{cases}\displaystyle\partial_{t}\left((\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})H_{t}\right)=-(\Delta_{\widetilde{E}}\otimes\mathrm{id})\left((\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})H_{t}\right),\\ \displaystyle\lim_{t\to 0}(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})H_{t}=0.\end{cases}$
By uniqueness of the solution to the heat equation, it follows that
$(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})H_{t}=0.$
The following proposition proves the existence of the propagator:
Proposition 2.8.
Let $(M,E,\omega)$ be a triple as in 2.4. Then
[…]
is a propagator.
Moreover, $P|_{M\times M\setminus\triangle}$ is smooth, where $\triangle:=\{(p,p)\in M\times M:p\in M\}$.
Proof.
The proof is similar to that of [4, Theorem 2.38], and we outline the main steps.
For $\epsilon,L\in(0,\infty)$, set
$).G_{\epsilon,L}:=\int_{\epsilon}^{L}dt\wedge H_{t}\in C^{\infty}\big{(}(0,\infty);\Gamma(\widetilde{E}\boxtimes\widetilde{E}$$\big{)}$.
Then by (2.5):
$\begin{cases}(\Delta_{\widetilde{E}}\otimes\mathrm{id})G_{\epsilon,L}=\int_{\epsilon}^{L}dt\wedge(\Delta_{\widetilde{E}}\otimes\mathrm{id})H_{t}=H_{\epsilon}-H_{L},\\ (\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})G_{\epsilon,L}=\int_{\epsilon}^{L}dt\wedge(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})H_{t}=0.\end{cases}$ (2.6)
[…]
$(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})\int_{\epsilon}^{L}dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}$
$=(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)G_{\epsilon,L}$
$=(\Delta_{\widetilde{E}}\otimes\mathrm{id})G_{\epsilon,L}=H_{\epsilon}-H_{L}.$
Moreover,
$\int_{\epsilon}^{L}dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}=\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)G_{\epsilon,L}\in\mathrm{Im}\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right).$
One can show that
$\lim_{\begin{subarray}{c}\epsilon\to 0\\ L\to\infty\end{subarray}}\int_{\epsilon}^{L}dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}$
exists (this is the nontrivial part of the argument), and since $\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}$ is a closed operator, we obtain:
$\begin{cases}(\bar{\partial}_{\widetilde{E}}\otimes\mathrm{id}+\mathrm{id}\otimes\bar{\partial}_{\widetilde{E}})P=\lim_{\epsilon\to 0}H_{\epsilon}-\lim_{L\to\infty}H_{L}=\delta-\mathcal{H},\\ P\in\mathrm{Im}(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}).\end{cases}$
Hence $P$ is a propagator.
To prove $P|_{M\times M\setminus\triangle}$ is smooth, observe:
$(\Delta_{\widetilde{E}}\otimes\mathrm{id})\lim_{\epsilon\to 0}G_{\epsilon,L}=\delta-H_{L}.$
Since the singularities of the right-hand side are supported on the diagonal $\triangle$, it follows from elliptic regularity (cf. [19]) that $\lim_{\epsilon\to 0}G_{\epsilon,L}$ is smooth off $\triangle$. Hence, so is
$\int_{0}^{L}dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}.$
By [4, Proposition 2.37], the integral
$\int_{L}^{\infty}dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}$
is smooth globally. Therefore,
$P=\int_{0}^{L}dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}+\int_{L}^{\infty}dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}$
is smooth away from the diagonal.
The mildness of the singularities of $P$ is crucial for defining Feynman graph integrals using the Cauchy principal value. In fact, the singularities of $P$ are fully determined by the singularities of the heat kernel $H_{t}$ at $t=0$:
Let $\epsilon\in(0,\infty)$, and choose a cutoff function $\eta\in C_{c}^{\infty}(\mathbb{R})$ such that $\eta(t)\equiv 1$ for $|t|\leq\epsilon$ and $\eta(t)=0$ for $|t|\geq 2\epsilon$.
Proposition 2.9.
The integral
$\int_{0}^{\infty}(1-\eta(t))\,dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}$
defines a smooth section over $M\times M$. In particular, the singularities of $P$ coincide with those of
$\int_{0}^{\infty}\eta(t)\,dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}.$
Proof.
This follows directly from [4, Proposition 2.37]. ∎
2.9 implies that to analyze the singularities of $P$, it suffices to understand the short-time asymptotics of the heat kernel $H_{t}$ near $t=0$.
We now introduce a useful notion:
Definition 2.10.
The propagator in Schwinger space is defined by
$),P_{t}:=-dt\wedge\big{(}\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\big{)}H_{t}+H_{t}\in\Omega^{*}\Big{(}(0,+\infty);\Gamma(\widetilde{E}\boxtimes\widetilde{E}$$\Big{)}$,
[…]
We observe that
$\int_{0}^{\infty}\eta(t)P_{t}=-\int_{0}^{\infty}\eta(t)\,dt\wedge\left(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\right)H_{t}.$
Hence, by 2.9, the singularities of $P$ are fully determined by the singularities of $P_{t}$ at $t=0$.
2.2 Feynman graph integrals
In this subsection, we introduce Feynman graph integrals for holomorphic field theories and establish their convergence in the sense of the Cauchy principal value. For the reader’s convenience, relevant definitions and properties of the Cauchy principal value are collected in Appendix B.
Let $\triangle$ denote the diagonal of $M\times M$. The blow-up of $M\times M$ along the diagonal is denoted by $\mathrm{Bl}_{\triangle}(M\times M)$, with the canonical projection map
$p:\mathrm{Bl}_{\triangle}(M\times M)\rightarrow M\times M,$
such that $p^{-1}(\triangle)$ is the exceptional divisor.
The propagator $P$ can be regarded as a bundle-valued smooth differential form on
$M\times M\setminus\triangle\cong\mathrm{Bl}_{\triangle}(M\times M)\setminus p^{-1}(\triangle).$
The theorem below shows that the propagator has singularities of the desired type.
Theorem 2.11.
Let $M$ be a closed real analytic Kähler manifold, $E\to M$ a real analytic Hermitian holomorphic vector bundle, and let
$\omega:E\otimes E\to K_{M}$
be a non-degenerate bundle map. Then:
  1. (1)
    $P$ has divisorial type singularities along $p^{-1}(\triangle)$ (see Definition B.12). Moreover, all holomorphic derivatives of $P$ also have divisorial type singularities along $p^{-1}(\triangle)$.
  2. (2)
    The pullback
    $\triangle^{*}P:=\lim_{\epsilon\rightarrow 0}\triangle^{*}\left(\int_{\epsilon}^{\infty}dt\wedge\big{(}\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}\big{)}H_{t}\right)$
    defines a smooth bundle-valued differential form on $M$, where $\triangle^{*}$ denotes the pullback of bundle-valued differential forms along the diagonal embedding
    $\triangle:M\rightarrow M\times M.$
    Moreover, the pullbacks of all holomorphic derivatives of $P$ along $\triangle$ are also smooth bundle-valued differential forms.
2.11 is the most technically involved result in this paper. The key ingredients for its proof are presented in section 3 and section 4, while the complete proof is given in section 4.2.1. In this section, we assume the validity of 2.11 and use it to construct Feynman graph integrals.
Remark 2.12.
The analyticity assumption in 2.11 is essential for our proof. It remains an interesting question whether this theorem holds without it.
Before introducing Feynman graph integrals, we recall the Fulton–MacPherson compactification of configuration spaces.
Definition 2.13.
Let $\vec{I}$ be an ordered finite set and $M$ a closed complex manifold. The configuration space of $M$ with respect to $\vec{I}$ is defined by
${\mathrm{Conf}}_{\vec{I}}(M)=\big{\{}(p_{1},\dots,p_{|\vec{I}|})\in M^{|\vec{I}|}:\,p_{i}\neq p_{j}\text{ for }i\neq j\big{\}},$
where $|\vec{I}|$ denotes the number of elements in $\vec{I}$.
For any subset $\vec{J}\subset\vec{I}$, define
$\triangle_{\vec{J}\subset\vec{I}}:=\big{\{}(p_{1},\dots,p_{|\vec{I}|})\in M^{|\vec{I}|}\,:\,p_{i}=p_{j}\text{ for all }i,j\in\vec{J}\,\big{\}}.$
Denote the blow-up of $M^{|\vec{I}|}$ along $\triangle_{\vec{J}\subset\vec{I}}$ by $\mathrm{Bl}_{\triangle_{\vec{J}\subset\vec{I}}}(M^{|\vec{I}|})$. We then have a natural map
$i:{\mathrm{Conf}}_{\vec{I}}(M)\rightarrow\prod_{\vec{J}\subset\vec{I}}\mathrm{Bl}_{\triangle_{\vec{J}\subset\vec{I}}}(M^{|\vec{I}|}).$
Definition 2.14.
The Fulton–MacPherson compactification of ${\mathrm{Conf}}_{\vec{I}}(M)$ is defined as the closure of $\mathrm{Im}(i)$ in $\prod_{\vec{J}\subset\vec{I}}\mathrm{Bl}_{\triangle_{\vec{J}\subset\vec{I}}}(M^{|\vec{I}|})$. We denote it by $\widetilde{{\mathrm{Conf}}}_{\vec{I}}(M)$.
Proposition 2.15.
Let $\vec{I}$ be an ordered finite set. Then $\widetilde{{\mathrm{Conf}}}_{\vec{I}}(M)$ is a closed complex manifold. Moreover, the complement $\widetilde{{\mathrm{Conf}}}_{\vec{I}}(M)\setminus{\mathrm{Conf}}_{\vec{I}}(M)$ is a simple normal crossing divisor.
Proof.
See [14]. ∎
Proposition 2.16.
Let $\vec{J}\subset\vec{I}$ be an ordered finite set. Then there exists a holomorphic map
$\pi_{\vec{J}\subset\vec{I}}:\widetilde{{\mathrm{Conf}}}_{\vec{I}}(M)\rightarrow\widetilde{{\mathrm{Conf}}}_{\vec{J}}(M)$
extending the natural forgetful map from ${\mathrm{Conf}}_{\vec{I}}(M)$ to ${\mathrm{Conf}}_{\vec{J}}(M)$. Moreover,
$\pi^{-1}_{\vec{J}\subset\vec{I}}\left(\widetilde{{\mathrm{Conf}}}_{\vec{J}}(M)\setminus{\mathrm{Conf}}_{\vec{J}}(M)\right)\subset\widetilde{{\mathrm{Conf}}}_{\vec{I}}(M)\setminus{\mathrm{Conf}}_{\vec{I}}(M).$
Proof.
See [14]. ∎
Now, let’s introduce the notion of Feynman graph integrals.
Let $\{e_{1},\ldots,e_{r}\}$ be a local holomorphic frame of the holomorphic vector bundle $E\to M$, where $r=\mathrm{rank}(E)$. Then any section $\varphi\in\Gamma(E)$ can be locally expressed as
$\varphi=\sum_{j=1}^{r}\varphi^{j}e_{j},$ (2.7)
where the coefficients $\varphi^{j}$ are smooth functions defined on some open subset of $M$.
Let $\mathbf{j}=(j_{1},\ldots,j_{n})\in\mathbb{Z}_{\geq 0}^{n}$. Given a local holomorphic coordinate system $(z_{1},\ldots,z_{n})$ on $M$, we define the holomorphic differential operator
$\partial^{\mathbf{j}}:=\frac{\partial^{|\mathbf{j}|}}{\partial z_{1}^{j_{1}}\cdots\partial z_{n}^{j_{n}}},\quad\text{where }|\mathbf{j}|:=j_{1}+\cdots+j_{n}.$ (2.8)
Definition 2.17.
A degree-$k$ holomorphic Lagrangian density $I$ is a linear map from $\Gamma(E)^{\otimes k}$ to $\Gamma(K_{M})$, which is locally of the form
$I(\varphi_{1},\ldots,\varphi_{k})=\sum c_{i_{1}\ldots i_{k}\,\mathbf{j}_{1}\ldots\mathbf{j}_{k}}\,d^{n}\mathbf{z}\,\wedge\partial^{\mathbf{j}_{1}}(\varphi_{1}^{i_{1}})\wedge\partial^{\mathbf{j}_{2}}(\varphi_{2}^{i_{2}})\wedge\cdots\wedge\partial^{\mathbf{j}_{k}}(\varphi_{k}^{i_{k}}),$ (2.9)
where each $\varphi_{l}$ is a local section of $E$ expressed as in (2.7), $\varphi_{l}^{i_{l}}$ denotes the $i_{l}$-th component function in this decomposition, $\partial^{\mathbf{j}_{l}}$ is a holomorphic differential operator of the form (2.8) with $\mathbf{j}_{l}\in\mathbb{Z}_{\geq 0}^{n}$, and the coefficients $c_{i_{1}\ldots i_{k}\,\mathbf{j}_{1}\ldots\mathbf{j}_{k}}$ are holomorphic functions. Lastly, $d^{n}\mathbf{z}=dz_{1}\wedge\cdots\wedge dz_{n}.$
The density $I$ naturally extends to $\Gamma(\widetilde{E}^{\otimes k})$ in a way compatible with (2.9), and we shall continue to denote it by $I$.
Remark 2.18.
For a coordinate-free and equivalent definition of holomorphic Lagrangian densities, see [36, Definition 2.15].
Definition 2.19.
A directed graph $\vec{\Gamma}$ consists of the following data:
  1. (a)
    An ordered set of vertices $\vec{\Gamma}_{0}$ and an ordered set of edges $\vec{\Gamma}_{1}$.
  2. (b)
    Two maps between sets
    $t,h:\vec{\Gamma}_{1}\rightarrow\vec{\Gamma}_{0}$
    which are the assignments of tail and head to each directed edge.
  3. (c)
    For each vertex $v\in\vec{\Gamma}_{0}$ of degree $\mathrm{deg}(v)\in\mathbb{Z}_{\geq 0}$, we assign a $\mathrm{deg}(v)$-degree holomorphic Lagrangian density $I_{v}$.
We will use $|\vec{\Gamma}_{0}|$ and $|\vec{\Gamma}_{1}|$ to denote the number of vertices and edges respectively.
Given a directed graph $\vec{\Gamma}$, we have
[…]
and
$\otimes_{v\in\vec{\Gamma}_{0}}I_{v}:\Gamma\Big{(}\prod_{v\in\vec{\Gamma}_{0}}M^{\deg(v)};\boxtimes_{v\in\vec{\Gamma}_{0}}\widetilde{E}^{\boxtimes\deg(v)}\Big{)}\to\Gamma\left(\boxtimes_{v\in\vec{\Gamma}_{0}}(K_{M}\otimes\Lambda^{\bullet}T_{0,1}^{*}M)\right).$
There exist natural isomorphisms
$\tau^{\vec{\Gamma}}:(M\times M)^{|\vec{\Gamma}_{1}|}\cong\prod_{v\in\vec{\Gamma}_{0}}M^{\deg(v)},$
and
$\tau^{\vec{\Gamma}}_{*}:\mathcal{D}\Big{(}(M\times M)^{|\vec{\Gamma}_{1}|};\boxtimes_{e\in\vec{\Gamma}_{1}}\widetilde{E}\boxtimes\widetilde{E}\Big{)}\cong\mathcal{D}\Big{(}\prod_{v\in\vec{\Gamma}_{0}}M^{\deg(v)};\boxtimes_{v\in\vec{\Gamma}_{0}}\widetilde{E}^{\boxtimes\deg(v)}\Big{)},$
which are determined by the structure of the directed graph $\vec{\Gamma}$. Hence, $\tau^{\vec{\Gamma}}_{*}\left(\otimes_{e\in\vec{\Gamma}_{1}}P\right)$ is an element of
$\mathcal{D}\Big{(}\prod_{v\in\vec{\Gamma}_{0}}M^{\deg(v)};\boxtimes_{v\in\vec{\Gamma}_{0}}\widetilde{E}^{\boxtimes\deg(v)}\Big{)},$
with singularities along the diagonals. Therefore, by Theorem 2.11,
$\left((\otimes_{v\in\vec{\Gamma}_{0}}I_{v})\circ\tau^{\vec{\Gamma}}_{*}\right)\left(\otimes_{e\in\vec{\Gamma}_{1}}P\right)$
is well-defined as a differential form on ${\mathrm{Conf}}_{\vec{\Gamma}_{0}}(M)$. We denote it by
$\left(\otimes_{v\in\vec{\Gamma}_{0}}I_{v},\otimes_{e\in\vec{\Gamma}_{1}}P\right)_{\vec{\Gamma}}\in\Gamma\left({\mathrm{Conf}}_{\vec{\Gamma}_{0}}(M);\boxtimes_{v\in\vec{\Gamma}_{0}}(K_{M}\otimes\Lambda^{\bullet}T_{0,1}^{*}M)\right).$
Theorem 2.20.
Let $(M,E,\omega)$ be a triple as in 2.11. Let $\vec{\Gamma}$ be a directed graph. Then
$\left(\otimes_{v\in\vec{\Gamma}_{0}}I_{v},\otimes_{e\in\vec{\Gamma}_{1}}P\right)_{\vec{\Gamma}}$
is a differential form on $\widetilde{{\mathrm{Conf}}}_{\vec{\Gamma}_{0}}(M)$ with divisorial type singularities along $\widetilde{{\mathrm{Conf}}}_{\vec{\Gamma}_{0}}(M)\setminus{\mathrm{Conf}}_{\vec{\Gamma}_{0}}(M)$ (see Definition B.12). In particular,
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{\widetilde{{\mathrm{Conf}}}_{\vec{\Gamma}_{0}}(M)}\left(\otimes_{v\in\vec{\Gamma}_{0}}I_{v},\otimes_{e\in\vec{\Gamma}_{1}}P\right)_{\vec{\Gamma}}$
is well-defined.
Proof.
By Theorem 2.11 and Proposition B.14, we conclude that
$\left(\otimes_{v\in\vec{\Gamma}_{0}}I_{v},\otimes_{e\in\vec{\Gamma}_{1}}P\right)_{\vec{\Gamma}}$
has divisorial type singularities. By the construction described in Appendix B, the integral is well-defined in the sense of Cauchy principal value. ∎
Definition 2.21.
Let $(M,E,\omega)$ be triples as in 2.11. Let $\vec{\Gamma}$ be a directed graph. The Feynman graph integral on $M$ is defined by:
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{\widetilde{{\mathrm{Conf}}}_{\vec{\Gamma}_{0}}(M)}\left(\otimes_{v\in\vec{\Gamma}_{0}}I_{v},\otimes_{e\in\vec{\Gamma}_{1}}P\right)_{\vec{\Gamma}}.$
Remark 2.22.
When $M$ is an elliptic curve and $E$ is the trivial vector bundle, the Feynman graph integrals can be computed explicitly, see [37].
In the remainder of this subsection, we introduce an alternative formulation of 2.20.
Definition 2.23.
Let $M$ be a closed real analytic Kähler manifold. A fake distance function $\tilde{\rho}:M\times M\rightarrow\mathbb{R}$ is a non-negative function satisfying:
  1. 1.
    $\tilde{\rho}^{2}$ is smooth and $\tilde{\rho}^{-1}(0)=\triangle$, where $\triangle:=\{(p,p)\in M\times M:p\in M\}$.
  2. 2.
    There exists an open neighborhood $U\subset M\times M$ of $\triangle$ such that
    $\tilde{\rho}^{2}|_{U}=\rho^{2}|_{U},$
    where $\rho$ is the distance function on $M$.
Lemma 2.24.
Let $M$ be a closed real analytic Kähler manifold, and let $\tilde{\rho}$ be a fake distance function. Then
$p^{*}\tilde{\rho}^{2}:\mathrm{Bl}_{\triangle}(M\times M)\rightarrow\mathbb{R}$
is a smooth defining function (see B.16) for the divisor $p^{-1}(\triangle)$, where $p:\mathrm{Bl}_{\triangle}(M\times M)\rightarrow M\times M$ is the canonical blow-up map along $\triangle:=\{(p,p)\in M\times M:p\in M\}$.
Proof.
Since being a smooth defining function is a local property, we may reduce to a local neighborhood in $\mathbb{C}^{n}$. It suffices to prove the statement over an open set $U\subset\mathbb{C}^{n}\times\mathbb{C}^{n}$ containing $(0,0)$, where we may assume $\tilde{\rho}=\rho$.
We identify $\mathrm{Bl}_{\triangle}(\mathbb{C}^{n}\times\mathbb{C}^{n})$ with the closure of the image of the natural embedding
$i:\mathbb{C}^{n}\times\mathbb{C}^{n}\setminus\triangle\hookrightarrow\left(\mathbb{C}^{n}\times\mathbb{C}^{n}\right)\times\left(\mathbb{C}^{n}\times\mathbb{C}^{n}\setminus\triangle\right)/\mathbb{C}^{*},\quad(\mathbf{z},\mathbf{w})\mapsto(\mathbf{z},\mathbf{w},\mathbf{z},\mathbf{w}),$
where the $\mathbb{C}^{*}$-action is given by $\lambda\cdot(\mathbf{z},\mathbf{w})=(\lambda\mathbf{z},\lambda\mathbf{w})$ for $\lambda\in\mathbb{C}^{*}$.
This yields an identification of $(\mathbb{C}^{n}\times\mathbb{C}^{n}\setminus\triangle)/\mathbb{C}^{*}$ with $\mathbb{C}^{n}\times\mathbb{P}^{n-1}$ via
$(\mathbf{z},\mathbf{w})\mapsto(\mathbf{w},\;[z_{1}-w_{1}:\cdots:z_{n}-w_{n}]).$
Under this identification, we can write
$\quad\mathrm{Bl}_{\triangle}(\mathbb{C}^{n}\times\mathbb{C}^{n})$
$\cong\{(\mathbf{z},\mathbf{w},[\lambda_{1}:\cdots:\lambda_{n}])\in\mathbb{C}^{n}\times\mathbb{C}^{n}\times\mathbb{P}^{n-1}:\mathbf{z}-\mathbf{w}=k\cdot\boldsymbol{\lambda}\text{ for some }k\in\mathbb{C}\},$
where $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{C}^{n}$.
Let $\{U_{i}\}_{i=1}^{n}$ be the standard affine cover of $\mathbb{P}^{n-1}$, corresponding to $\lambda_{i}\neq 0$. Over $U_{1}$, we can use local coordinates
$(z_{1}-w_{1},\lambda_{2},\ldots,\lambda_{n},w_{1},\ldots,w_{n}).$
Using the expansion from (A.2), we compute:
$\frac{\rho^{2}(\mathbf{z},\bar{\mathbf{z}},\mathbf{w},\bar{\mathbf{w}})}{|z_{1}-w_{1}|^{2}}$ $=1+\sum_{i=2}^{n}\lambda_{i}\bar{\lambda}_{i}$ (2.10)
$\quad+\Big{(}f_{1\bar{1}}+\sum_{i=2}^{n}\lambda_{i}f_{i\bar{1}}+\sum_{i=2}^{n}\bar{\lambda}_{i}f_{1\bar{i}}+\sum_{i,j=2}^{n}\lambda_{i}\bar{\lambda}_{j}f_{i\bar{j}}\Big{)},$ (2.11)
where $f_{i\bar{j}}(0,0,0,0)=0$.
By shrinking $U$ if necessary, we may ensure $|f_{i\bar{j}}|<\min\left\{\frac{1}{3},\frac{1}{3(n-1)},\frac{1}{3(n-1)^{2}}\right\}$. Then the absolute value of the term (2.10) dominates that of (2.11), so the full expression is a positive smooth function on $p^{-1}(U)\cap U_{1}$. Thus, $\rho^{2}$ is a smooth defining function on this chart. A similar argument works on other charts $U_{i}$, completing the proof. ∎
Theorem 2.25.
Let $(M,E,\omega)$ be triples as in 2.11. Let $\vec{\Gamma}$ be a directed graph, and $\tilde{\rho}$ a fake distance function. Then we have
$\lim_{\epsilon\rightarrow 0}\int_{\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}_{ij}>\epsilon}\left(\otimes_{v\in\vec{\Gamma}_{0}}I_{v},\otimes_{e\in\vec{\Gamma}_{1}}P\right)_{\vec{\Gamma}}=\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{\widetilde{{\mathrm{Conf}}}_{\vec{\Gamma}_{0}}(M)}\left(\otimes_{v\in\vec{\Gamma}_{0}}I_{v},\otimes_{e\in\vec{\Gamma}_{1}}P\right)_{\vec{\Gamma}},$
where $\tilde{\rho}^{2}_{ij}=\pi_{\{i<j\}\subset\vec{\Gamma}_{0}}^{*}(\tilde{\rho}^{2})$.
Proof.
Consider the following natural map:
$\prod_{i<j\in\vec{\Gamma}_{0}}\pi_{\{i<j\}\subset\vec{\Gamma}_{0}}:\widetilde{{\mathrm{Conf}}}_{\vec{\Gamma}_{0}}(M)\rightarrow\prod_{i<j\in\vec{\Gamma}_{0}}\widetilde{{\mathrm{Conf}}}_{\{i<j\}}(M).$
We note that $\widetilde{{\mathrm{Conf}}}_{\{i<j\}}(M)\cong\mathrm{Bl}_{\triangle}(M\times M)$, so the function $\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}_{ij}$ is a smooth defining function of the divisor
$\bigcup_{i<j\in\vec{\Gamma}_{0}}p_{ij}^{-1}(\triangle)\subset\prod_{i<j\in\vec{\Gamma}_{0}}\widetilde{{\mathrm{Conf}}}_{\{i<j\}}(M),$
where $p_{ij}$ is the composition
$\prod_{i^{\prime}<j^{\prime}\in\vec{\Gamma}_{0}}\widetilde{{\mathrm{Conf}}}_{\{i^{\prime}<j^{\prime}\}}(M)\rightarrow\widetilde{{\mathrm{Conf}}}_{\{i<j\}}(M)\rightarrow M\times M.$
Since
$\Big{(}\prod_{i<j\in\vec{\Gamma}_{0}}\pi_{\{i<j\}\subset\vec{\Gamma}_{0}}\Big{)}^{-1}\Big{(}\bigcup_{i<j\in\vec{\Gamma}_{0}}p_{ij}^{-1}(\triangle)\Big{)}=\widetilde{{\mathrm{Conf}}}_{\vec{\Gamma}_{0}}(M)\setminus{\mathrm{Conf}}_{\vec{\Gamma}_{0}}(M),$
it follows from Proposition B.17 that $\prod_{i<j\in\vec{\Gamma}_{0}}\tilde{\rho}^{2}_{ij}$ is a smooth defining function of
$\widetilde{{\mathrm{Conf}}}_{\vec{\Gamma}_{0}}(M)\setminus{\mathrm{Conf}}_{\vec{\Gamma}_{0}}(M).$
The theorem then follows from Proposition B.18. ∎
Remark 2.26.
Theorem 2.25 provides a definition of Feynman graph integrals that does not rely on the Fulton–MacPherson compactification. This shows that the notion of Feynman graph integral introduced in this paper is intrinsic.
3 Regular Expressions
In this section, we introduce a structure we call regular expressions, which serves as an analog of divisorial type singularities in the context of heat kernels. More precisely, in Section 3.1, we define regular expressions and establish their basic properties. In Section 3.2, we show the connection between regular expressions and divisorial type singularities.
3.1 Regular expressions and its filtration
Definition 3.1.
Let $U\subset\mathbb{C}^{n}$ be a open subset, $E$ is a trivial holomorphic vector bundle on $\mathbb{C}^{n}$, and $\widetilde{E}=E\otimes\Lambda^{\bullet}T_{0,1}^{*}\mathbb{C}^{n}$. The linear space of regular expressions $S_{{\mathrm{reg}}}(U\times U)\subset\Omega^{*}\Big{(}(0,\infty);\Gamma\big{(}(\widetilde{E}\boxtimes\widetilde{E})|_{U\times U}\big{)}\Big{)}$ is generated by expressions of the form
$a(\mathbf{y},d\mathbf{y})\,b,$
where $a(\mathbf{y},d\mathbf{y})$ is a polynomial in the variables
$\mathbf{y}=(y_{1},y_{2},\ldots,y_{n}),\quad\text{and}\quad d\mathbf{y}=(dy_{1},dy_{2},\ldots,dy_{n}),$
with
$y_{i}=\frac{\overline{z_{i}}-\overline{w_{i}}}{t}.$
Here, $d$ denotes the de Rham differential on $(0,\infty)\times U\times U$, so
$dy_{i}=d\left(\frac{\overline{z_{i}}-\overline{w_{i}}}{t}\right)=\frac{d(\overline{z_{i}}-\overline{w_{i}})}{t}-\frac{(\overline{z_{i}}-\overline{w_{i}})\,dt}{t^{2}}.$
The factor $b\in\Omega^{*}\Big{(}(0,\infty);\Gamma\big{(}(\widetilde{E}\boxtimes\widetilde{E})|_{U\times U}\big{)}\Big{)}$ is assumed to extend smoothly to a differential form in $\Omega^{*}\Big{(}[0,+\infty);\Gamma\big{(}(\widetilde{E}\boxtimes\widetilde{E})|_{U\times U}\big{)}\Big{)}$.
Proposition 3.2.
Let $f:U\subset\mathbb{C}^{n}\rightarrow W\subset\mathbb{C}^{n}$ be a biholomorphism, then $f^{*}(S_{{\mathrm{reg}}}(W\times W))=S_{{\mathrm{reg}}}(U\times U)$.
Proof.
Let $(\mathbf{z},\mathbf{w})\in U\times U$, we have
$f_{i}(\mathbf{z})=f_{i}(\mathbf{w})+\sum_{j=1}^{n}f_{ij}(\mathbf{z},\mathbf{w})(z_{j}-w_{j}),$
for some holomorphic functions $f_{ij}$.
Now let $y_{i}^{\prime}:=\frac{\bar{f}_{i}(\mathbf{z})-\bar{f}_{i}(\mathbf{w})}{t}$. Then
$y_{i}^{\prime}=\frac{\bar{f}_{i}(\mathbf{z})-\bar{f}_{i}(\mathbf{w})}{t}=\sum_{j}\overline{f_{ij}(\mathbf{z},\mathbf{w})}\cdot\frac{\bar{z}_{j}-\bar{w}_{j}}{t}=\sum_{j}\overline{f_{ij}(\mathbf{z},\mathbf{w})}\cdot y_{j}\in S_{{\mathrm{reg}}}(U\times U)$
and
$dy_{i}^{\prime}=\sum_{j}\left(d\overline{f_{ij}(\mathbf{z},\mathbf{w})}\right)y_{j}+\overline{f_{ij}(\mathbf{z},\mathbf{w})}\,dy_{j}\in S_{{\mathrm{reg}}}(U\times U),$
where we recall that $d$ denotes the de Rham differential on $(0,\infty)\times M\times M$. This proves our claim. ∎
Remark 3.3.
Proposition 3.2 shows that the definition of regular expressions is independent of the choice of coordinates. This allows us to define the concept globally on an complex manifold, see 4.1.
In general, it is difficult to check whether a given element in $)\Omega^{*}\big{(}(0,\infty);\Gamma(\widetilde{E}\boxtimes\widetilde{E}$$\big{)}$ is a regular expression. To obtain more characterizations of regular expressions, we restrict ourselves to analytic expressions.
Definition 3.4.
Let $U$ be an open subset of $\mathbb{C}^{n}$. For any $r>0$, we define
$N(U,r)=\Big{\{}(\mathbf{z},\mathbf{w})\in\mathbb{C}^{n}\times\mathbb{C}^{n}:\mathbf{w}\in U,|\mathbf{z}-\mathbf{w}|<r\Big{\}}.$
On $N(U,r)$, we introduce the coordinates
$(\tilde{\mathbf{z}},\mathbf{w})=(\mathbf{z}-\mathbf{w},\mathbf{w}).$
We will always use these coordinates for $N(U,r)$ in this paper.
Definition 3.5.
Let $p\in\mathbb{C}^{n}$, $U\subset\mathbb{C}^{n}$ be a neighborhood of $p$, $E$ is a trivial holomorphic vector bundle on $\mathbb{C}^{n}$, and $\widetilde{E}=E\otimes\Lambda^{\bullet}T_{0,1}^{*}\mathbb{C}^{n}$. For $r>0$, the linear space of analytic expressions $A\big{(}N(U,r);p\big{)}\subset\Omega^{*}\big{(}(0,\infty);\Gamma((\widetilde{E}\boxtimes\widetilde{E})|_{N(U,r)})\big{)}$ is generated by expressions of the form
$a\,t^{m}+a^{\prime}t^{m^{\prime}}dt,$
where $m,m^{\prime}\in\mathbb{Z}$, and $a,a^{\prime}$ are convergent power series at $(p,p)$ valued in $(\widetilde{E}\boxtimes\widetilde{E})|_{N(U,r)}$. Similarly, the linear space of regular analytic expressions
$A_{{\mathrm{reg}}}(N(U,r);p)\subset\Omega^{*}((0,\infty);\Gamma((\widetilde{E}\boxtimes\widetilde{E}|_{N(U,r)})))$
is generated by expressions of the form
$a(\mathbf{y},d\mathbf{y})\,b\,t^{m}+a^{\prime}(\mathbf{y},d\mathbf{y})\,b^{\prime}\,t^{m^{\prime}}dt,$
where $m,m^{\prime}\in\mathbb{Z}_{\geq 0}$, $a(\mathbf{y},d\mathbf{y}),a^{\prime}(\mathbf{y},d\mathbf{y})$ are polynomials in the variables
$\mathbf{y}=(y_{1},y_{2},\ldots,y_{n}),\quad\text{and}\quad d\mathbf{y}=(dy_{1},dy_{2},\ldots,dy_{n}),$
with
$y_{i}=\frac{\bar{\tilde{z}}_{i}}{t}=\frac{\overline{z_{i}}-\overline{w_{i}}}{t}.$
The factors $b,b^{\prime}$ are convergent power series at $(p,p)$ of $(\widetilde{E}\boxtimes\widetilde{E})|_{N(U,r)}$.
Given $A(N(U,r);p)$, where $p\in U\subset\mathbb{C}^{n}$ and $r>0$, we will define a scaling action—namely, a version of Getzler’s rescaling—on $A(N(U,r);p)$. For $\epsilon\in(0,1]$, the scaling operator $\delta_{\epsilon}$ is defined by
$\delta_{\epsilon}\big{(}a(\tilde{{\mathbf{z}}},{\bar{\tilde{{\mathbf{z}}}}},t,{\mathbf{w}},\bar{{\mathbf{w}}},d{\bar{\tilde{{\mathbf{z}}}}},d\bar{{\mathbf{w}}},dt)\big{)}=a(\tilde{{\mathbf{z}}},\epsilon{\bar{\tilde{{\mathbf{z}}}}},\epsilon t,{\mathbf{w}},\bar{{\mathbf{w}}},\epsilon d{\bar{\tilde{{\mathbf{z}}}}},d\bar{{\mathbf{w}}},\epsilon dt),$ (3.1)
for any $a\in A\big{(}N(U,r);p\big{)}$.
This scaling action defines a filtration on $A(N(U,r);p)$:
Definition 3.6.
For $m\in\mathbb{Z}$, we define
$F_{m}A\big{(}N(U,r);p\big{)}=\Big{\{}a\in A\big{(}N(U,r);p\big{)}\;:\;\lim_{\epsilon\rightarrow 0}\epsilon^{-m}\delta_{\epsilon}(a)\text{ exists}\Big{\}}.$
This gives a descending filtration:
$\cdots\supset F_{m}A\big{(}N(U,r);p\big{)}\supset F_{m+1}A\big{(}N(U,r);p\big{)}\supset\cdots.$
We now state two propositions regarding the filtration $F_{*}A\big{(}N(U,r);p\big{)}$.
Proposition 3.7.
Let $a\in A\big{(}N(U,r);p\big{)}$. Then $a\in F_{m}A\big{(}N(U,r);p\big{)}$ if and only if
$\lim_{\epsilon\rightarrow 0}\epsilon^{-m+1}\delta_{\epsilon}(a)=0.$
Proof.
This follows from the fact that $\delta_{\epsilon}a$ admits a real Laurent expansion in $\epsilon$ of the form
$\delta_{\epsilon}a=\sum_{i=-l}^{\infty}a_{i}\epsilon^{i},$
for some $l\in\mathbb{Z}_{\geq 0}$, with $a_{i}\in A\big{(}N(U,r);p\big{)}$. The limit $\epsilon^{-m}\delta_{\epsilon}(a)$ exists if and only if all terms with $i<m$ vanish, which is equivalent to saying $\epsilon^{-m+1}\delta_{\epsilon}(a)\to 0$ as $\epsilon\to 0$. ∎
Proposition 3.8.
Let $f:U\subset\mathbb{C}^{n}\rightarrow V\subset\mathbb{C}^{n}$ be a holomorphic map with $p\in U$ and $r,r^{\prime}>0$. Suppose that $f\big{(}N(U,r)\big{)}\subset N(V,r^{\prime})$ and that
$f^{*}\Big{(}A\big{(}N(V,r^{\prime});f(p)\big{)}\Big{)}\subset A\big{(}N(U,r);p\big{)}.$
Then for any $m\in\mathbb{Z}$, we have
$f^{*}\big{(}F_{m}A\big{(}N(V,r^{\prime});f(p)\big{)}\big{)}\subset F_{m}A\big{(}N(U,r);p\big{)}.$
Proof.
Let $(\tilde{{\mathbf{z}}},{\mathbf{w}})$ and $(\tilde{{\mathbf{z}}}^{\prime},{\mathbf{w}}^{\prime})$ denote the coordinates on $N(U,r)$ and $N(V,r^{\prime})$, respectively. For $\epsilon>0$, denote by $\delta_{\epsilon}$ and $\delta^{\prime}_{\epsilon}$ the scaling operators on $N(U,r)$ and $N(V,r^{\prime})$. Then
$\begin{cases}\delta^{\prime}_{\epsilon}\bar{\tilde{z}}^{\prime}_{i}=\epsilon\big{(}\bar{f}_{i}({\mathbf{z}})-\bar{f}_{i}({\mathbf{w}})\big{)}=\delta_{\epsilon}\bar{\tilde{z}}^{\prime}_{i}+O(\epsilon),\\ \delta^{\prime}_{\epsilon}t=\delta_{\epsilon}t=\epsilon t.\end{cases}$
For $a^{\prime}\in A(N(V,r^{\prime});f(p))$, we can write
$a^{\prime}=\sum_{i=l_{1}}^{l_{2}}t^{i}\Big{(}\sum_{\begin{subarray}{c}\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}\\ \mathbf{k}\in\{0,1\}^{n}\end{subarray}}b^{i}_{\mathbf{j}\mathbf{k}}\bar{\tilde{{\mathbf{z}}}}^{\prime\mathbf{j}}d\bar{\tilde{{\mathbf{z}}}}^{\prime\mathbf{k}}+\sum_{\begin{subarray}{c}\mathbf{j}\in\mathbb{Z}_{\geq 0}^{n}\\ \mathbf{k}\in\{0,1\}^{n}\end{subarray}}c^{i}_{\mathbf{j}\mathbf{k}}\bar{\tilde{{\mathbf{z}}}}^{\prime\mathbf{j}}d\bar{\tilde{{\mathbf{z}}}}^{\prime\mathbf{k}}dt\Big{)},$
where $b^{i}_{\mathbf{j}\mathbf{k}}$ and $c^{i}_{\mathbf{j}\mathbf{k}}$ are analytic sections of $(\widetilde{E}\boxtimes\widetilde{E})|_{N(V,r^{\prime})}$ that are independent of $\bar{\tilde{{\mathbf{z}}}}$ and $d\bar{\tilde{{\mathbf{z}}}}$. Here for ${\mathbf{i}}\in\mathbb{Z}^{n}_{\geq 0}$, ${\mathbf{z}}^{\mathbf{i}}:=z_{1}^{i_{1}}\cdots z_{n}^{i_{n}},$ and for ${\mathbf{i}}=(i_{1},\cdots,i_{n})\in\{0,1\}^{n}$ such that $i_{j_{1}}=\cdots i_{j_{l}}=1$, $dz^{\mathbf{i}}:=dz_{j_{1}}\wedge\cdots\wedge dz_{j_{l}}.$
If $a^{\prime}\in F_{m}A(N(V,r^{\prime});f(p))$, then
$\begin{cases}b^{i}_{\mathbf{j}\mathbf{k}}=0&\text{if }|\mathbf{j}|+|\mathbf{k}|+i<m,\\ c^{i}_{\mathbf{j}\mathbf{k}}=0&\text{if }|\mathbf{j}|+|\mathbf{k}|+i+1<m.\end{cases}$
Here for ${\mathbf{j}}=(j_{1},\cdots,j_{n})\in\mathbb{Z}_{\geq 0}^{n}\text{ or }\{0,1\}^{n},|{\mathbf{j}}|:=j_{1}+\cdots+j_{n}$.
Then we have
$\epsilon^{-m+1}\delta_{\epsilon}f^{*}(a^{\prime})=f^{*}(\epsilon^{-m+1}\delta^{\prime}_{\epsilon}a^{\prime})+O(\epsilon).$
By Proposition 3.7, we conclude that $f^{*}(a^{\prime})\in F_{m}A(N(U,r);p)$. ∎
Remark 3.9.
Proposition 3.8 shows that the filtration $F_{*}A(N(U,r);p)$ is independent of the choice of coordinates. This allows us to define the filtration globally on complex manifold. The idea of introducing scaling actions and filtrations is motivated by Getzler’s rescaling, which plays a crucial role in the study of heat kernel expansions and index theorems. See [4, 15] for further details.
Consider the vector field
$X:=t\partial_{t}+\sum_{i=1}^{n}\bar{\tilde{z}}_{i}\partial_{\bar{\tilde{z}}_{i}}.$ (3.2)
Using the filtration $F_{*}A(N(U,r);p)$, we have the following characterization of analytic regular expressions:
Proposition 3.10.
Let $p\in U\subset\mathbb{C}^{n}$ and $r>0$. An analytic expression $a\in A(N(U,r);p)$ lies in $A_{\mathrm{reg}}(N(U,r);p)$ if and only if $a\in F_{0}A(N(U,r);p)$ and
$\iota_{X}\left(\lim_{\epsilon\to 0}\delta_{\epsilon}(a)\right)=0,$
where $\iota_{X}$ denotes the interior product with respect to the vector field $X$.
Proof.
We first prove the “only if” direction. Suppose $a\in A_{\mathrm{reg}}(N(U,r);p)$. Then
$\lim_{\epsilon\to 0}\delta_{\epsilon}(a)=\sum_{k=0}^{m}a_{k}(\mathbf{y},d\mathbf{y})\cdot b_{k},$
for some polynomials $a_{k}(\mathbf{y},d\mathbf{y})$ in the variables
$\mathbf{y}=(y_{1},\dots,y_{n}),\quad d\mathbf{y}=(dy_{1},\dots,dy_{n}),\quad y_{j}=\frac{\bar{\tilde{z}}_{j}}{t}=\frac{\bar{z}_{j}-\bar{w}_{j}}{t},$
and some $b_{k}$, which are analytic sections of $\widetilde{E}\boxtimes\widetilde{E}$ over $N(U,r)$, independent of $\bar{\tilde{{\mathbf{z}}}}$ and $d\bar{\tilde{{\mathbf{z}}}}$.
Since $\iota_{X}(y_{j})=\iota_{X}(dy_{j})=\iota_{X}(b_{k})=0$, it follows that
$\iota_{X}\big{(}\lim_{\epsilon\to 0}\delta_{\epsilon}(a)\big{)}=0.$
Moreover, by definition of regular expressions, $a\in F_{0}A(N(U,r);p)$.
For the “if” direction, suppose $a\in F_{0}A(N(U,r);p)$ and $\iota_{X}\left(\lim_{\epsilon\to 0}\delta_{\epsilon}(a)\right)=0$. Then we can write
$\lim_{\epsilon\to 0}\delta_{\epsilon}(a)=\sum_{k=0}^{m}\left(a_{k}(\mathbf{y},d\mathbf{y})\cdot b_{k}+\frac{dt}{t}\cdot a_{k}^{\prime}(\mathbf{y},d\mathbf{y})\cdot b_{k}^{\prime}\right),$
for some polynomials $a_{k}$, $a_{k}^{\prime}$ in the variables $\mathbf{y},d\mathbf{y}$, and some analytic sections $b_{k}$, $b_{k}^{\prime}$ of $\widetilde{E}\boxtimes\widetilde{E}$ over $N(U,r)$, which are independent of $\bar{\tilde{{\mathbf{z}}}}$ and $d\bar{\tilde{{\mathbf{z}}}}$. By assumption,
$\iota_{X}\big{(}\lim_{\epsilon\to 0}\delta_{\epsilon}(a)\big{)}=\sum_{k=0}^{m}a_{k}^{\prime}(\mathbf{y},d\mathbf{y})\cdot b_{k}^{\prime}=0.$
Since
$\sum_{k=0}^{m}a_{k}(\mathbf{y},d\mathbf{y})\cdot b_{k}\in A_{{\mathrm{reg}}}\big{(}N(U,r);p\big{)},$
and it is easily checked that
$a-\lim_{\epsilon\to 0}\delta_{\epsilon}(a)\in A_{{\mathrm{reg}}}\big{(}N(U,r);p\big{)},$
we conclude that $a\in A_{{\mathrm{reg}}}\big{(}N(U,r);p\big{)}$.
We now present an important fiberwise criterion for regularity.
Let $\pi_{2}:N(U,r)\subset\mathbb{C}^{n}\times\mathbb{C}^{n}\to\mathbb{C}^{n}$ denote the projection to the second factor:
$\pi_{2}(\tilde{{\mathbf{z}}},{\mathbf{w}})={\mathbf{w}}.$
Definition 3.11.
Let $p\in U\subset\mathbb{C}^{n}$ and $r>0$. An analytic expression $a\in A(N(U,r);p)$ is said to be regular along $\pi_{2}^{-1}(p^{\prime})$ for $p^{\prime}\in U$ if there exists $a^{\prime}\in A_{\mathrm{reg}}(N(U,r);p)$ such that
$a|_{\pi_{2}^{-1}(p^{\prime})}=a^{\prime}|_{\pi_{2}^{-1}(p^{\prime})}.$
We write
$N(p^{\prime},r):=\{{\mathbf{z}}\in\mathbb{C}^{n}:|{\mathbf{z}}-p^{\prime}|<r\}$
and use $A_{\mathrm{reg}}(N(p^{\prime},r);p)$ to denote the space of analytic expressions that are regular along $\pi_{2}^{-1}(p^{\prime})\cong N(p^{\prime},r)$.
Proposition 3.12.
Let $p\in U\subset\mathbb{C}^{n}$ and $r>0$. Then $a\in A_{\mathrm{reg}}(N(U,r);p)$ if and only if $a\in A_{\mathrm{reg}}(N(p^{\prime},r);p)$ for every $p^{\prime}\in U$.
Proof.
By Proposition 3.10, $a$ is regular if and only if
$\begin{cases}\lim_{\epsilon\to 0}\epsilon\delta_{\epsilon}(a)=0,\\ \iota_{X}\left(\lim_{\epsilon\to 0}\delta_{\epsilon}(a)\right)=0.\end{cases}$
Since the vector field $\sum\bar{\tilde{z}}_{i}\partial_{\bar{\tilde{z}}_{i}}$ is tangent to the fibers of $\pi_{2}$, both conditions are equivalent to holding along every fiber $\pi_{2}^{-1}(p^{\prime})$. This proves the proposition. ∎
Remark 3.13.
Proposition 3.12 reduces the verification of regularity to a fiberwise check. As we will see in later sections, this greatly simplifies the proof of the regularity of the propagator.
3.2 Regular expressions and divisorial type singularities
In this subsection, we establish the connection between regular expressions of and divisorial type singularities. Let $U\subset\mathbb{C}^{n}$ be an open subset, and let $E$ be a trivial holomorphic vector bundle over $\mathbb{C}^{n}$. Set $\widetilde{E}=E\otimes\Lambda^{\bullet}T^{*}_{0,1}\mathbb{C}^{n}$. We further assume that $U\subset\mathbb{C}^{n}$ is equipped with a real analytic Kähler metric such that the squared distance function $\rho^{2}:U\times U\to\mathbb{R}$ is smooth.
Let $\eta\in C_{c}^{\infty}(\mathbb{R})$ be a cutoff function satisfying $\eta(t)\equiv 1$ for $|t|\leq 1$, and $\eta(t)=0$ for $|t|\geq 2$.
We begin with the following:
Proposition 3.14.
Let $C\in S_{\mathrm{reg}}(U\times U)$. Then there exist polynomials $\alpha_{k}(\tilde{y}_{i},d\tilde{y}_{i})$, $\tilde{\alpha}_{k}(\tilde{y}_{i},d\tilde{y}_{i})$ and smooth differential forms $\beta_{k},\tilde{\beta}_{k}\in\Gamma(\widetilde{E}\boxtimes\widetilde{E}),k=1,\cdots,m$ such that
$\int_{0}^{\infty}\eta(t)\,e^{-\frac{\rho^{2}}{2t}}\,C=\sum_{k=0}^{m}\alpha_{k}(\tilde{y}_{i},d\tilde{y}_{i})\,\beta_{k}+\sum_{k=0}^{m}\tilde{\alpha}_{k}(\tilde{y}_{i},d\tilde{y}_{i})\,\tilde{\beta}_{k}\ln(\rho^{2}),$ (3.3)
where
$\tilde{y}_{i}=\frac{\bar{z}_{i}-\bar{w}_{i}}{\rho^{2}({\mathbf{z}},\bar{{\mathbf{z}}},{\mathbf{w}},\bar{{\mathbf{w}}})},$
and the integral is taken over the $t$-variable. That is, if
$),w=\alpha+dt\wedge\beta,\quad\text{with }\alpha,\beta\in C^{\infty}\big{(}(0,\infty);\Gamma(\widetilde{E}\boxtimes\widetilde{E}$$\big{)}$,
then
$\int_{0}^{\infty}w:=\int_{0}^{\infty}\beta\,dt.$
Proof.
It suffices to consider the case where
$\eta C_{1}=a({\mathbf{y}},d{\mathbf{y}})b_{1},$ (3.4)
or
$\eta C_{2}=a({\mathbf{y}},d{\mathbf{y}})b_{2}dt,$ (3.5)
where $a$ is a monomial of degree $m$ with $m\geq 0$, and $)b_{1},b_{2}\in C_{c}^{\infty}([0,\infty),\Gamma(\widetilde{E}\boxtimes\widetilde{E}$)).
Let $\hat{b}_{i}(s)$ denote the Mellin transform of $b_{i}(t)$, for $i=1,2$:
$\hat{b}_{i}(s)=\int_{0}^{\infty}\tau^{s-1}b_{i}(\tau)\,d\tau.$ (3.6)
[…]
By the inverse Mellin transform, we obtain
$b_{i}(t)=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}t^{-s}\hat{b}_{i}(s)ds,$ (3.7)
for some constant $c>0$.
Applying Fubini’s theorem, (3.6) and (3.7), we derive, for $c>0$,
$\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C_{1}$ $=\frac{1}{2\pi i}\int_{0}^{\infty}e^{-\frac{\rho^{2}}{2t}}a({\mathbf{y}},d{\mathbf{y}})\int_{c-i\infty}^{c+i\infty}t^{-s}\hat{b}_{1}(s)ds$ (3.8)
$=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}\left(\int_{0}^{\infty}e^{-\frac{\rho^{2}}{2t}}a({\mathbf{y}},d{\mathbf{y}})t^{-s}\right)\hat{b}_{1}(s)ds.$
Next, we evaluate the inner integral. Changing variables via $t^{\prime}=\frac{\rho^{2}}{2t}$, we rewrite it as
$\int_{0}^{\infty}e^{-\frac{\rho^{2}}{2t}}a({\mathbf{y}},d{\mathbf{y}})t^{-s}=\left(\frac{\rho}{\sqrt{2}}\right)^{-2s}\int_{0}^{\infty}e^{-t^{\prime}}a\big{(}t^{\prime}\tilde{{\mathbf{y}}},d(t^{\prime}\tilde{{\mathbf{y}}})\big{)}(t^{\prime})^{s}.$ (3.9)
Expanding $)a(t^{\prime}\tilde{{\mathbf{y}}},d(t^{\prime}\tilde{{\mathbf{y}}}$)) in powers of $t^{\prime}$, we obtain
$a\big{(}t^{\prime}\tilde{{\mathbf{y}}},d(t^{\prime}\tilde{{\mathbf{y}}})\big{)}=a_{1}(\tilde{{\mathbf{y}}},d\tilde{{\mathbf{y}}})(t^{\prime})^{m}+a_{2}(\tilde{{\mathbf{y}}},d\tilde{{\mathbf{y}}})(t^{\prime})^{m-1}dt^{\prime},$ (3.10)
for some polynomials $a_{1},a_{2}$.
Substituting (3.10) into (3.9), we obtain
$\int_{0}^{\infty}e^{-\frac{\rho^{2}}{2t}}a({\mathbf{y}},d{\mathbf{y}})t^{-s}=\left(\frac{\rho}{\sqrt{2}}\right)^{-2s}a_{2}(\tilde{{\mathbf{y}}},d\tilde{{\mathbf{y}}})\Gamma(s+m).$ (3.11)
Combining (3.8) and (3.11), we deduce that for $c>0$,
$\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C_{1}=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}a_{2}(\tilde{{\mathbf{y}}},d\tilde{{\mathbf{y}}})\hat{b}_{1}(s)\left(\frac{\rho}{\sqrt{2}}\right)^{-2s}\Gamma(s+m)ds.$ (3.12)
Similarly, for $c>1$, we find
$\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C_{2}=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}a_{1}(\tilde{{\mathbf{y}}},d\tilde{{\mathbf{y}}})\hat{b}_{2}(s)\left(\frac{\rho}{\sqrt{2}}\right)^{2-2s}\Gamma(s+m-1)ds.$ (3.13)
Since $\Gamma(s)$ has simple poles at $s=0,-1,-2,\dots$, it follows that $\hat{b}_{1}(s)\Gamma(s+m)$ has at most double poles at these points, while $\hat{b}_{2}(s)\Gamma(s+m-1)$ has at most double poles at $s=0,-1,-2,\dots$, and a simple pole at $s=1$.
From (3.19) in 3.15, 3.16, along with (3.12) and (3.13), we conclude that (3.3) holds.
3.15 and 3.16 are instrumental in the proof of 3.14. Although they follow from the standard theory of the Mellin transform, we could not find a reference that covers the precise versions we require, so we include a proof here for completeness.
Lemma 3.15.
Let $c>1$ be a real number, and let $h(s)$ be a meromorphic function with at most double poles at $s=0,-1,-2,\dots$, and a simple pole at $s=1$.
Furthermore, assume:
  1. (a)
    Near s=ks=-k for each $k=0,1,2,\dots$, the function $h(s)$ has the Laurent series expansion
    $h(s)=\frac{a_{k}}{(s+k)^{2}}+\frac{b_{k}}{s+k}+\cdots.$ (3.14)
    Near $s=1$, $h(s)$ has the Laurent series expansion
    $h(s)=\frac{b_{-1}}{s-1}+\cdots.$ (3.15)
  2. (b)
    For any $N>2$, consider the region
    $S_{-N,c}:=\big{\{}s\in\mathbb{C}:{\mathrm{Re}}(s)\in[-N,c],\ |\mathrm{Im}(s)|\geq 1\big{\}}.$
    For any integer $l\geq 1$, there exists a constant $C_{N,l}>0$ such that for all $s\in S_{-N,c}$,
    $|h(s)|\leq\frac{C_{N,l}}{|\mathrm{Im}(s)|^{l}}.$ (3.16)
  3. (c)
    For any integer $t\in\mathbb{R}$, let
    $L_{t}:=\{s\in\mathbb{C}:{\mathrm{Re}}(s)=t\}.$
    Then for each integer $k\geq 2$, there exists a constant $M_{k}>0$ such that for all $s\in L_{-k-\frac{1}{2}}$,
    $|h(s)|\leq M_{k}.$ (3.17)
Then the function $g(u)$ defined by
$g(u):=\int_{L_{c}}h(s)\,u^{-s}\,ds$
is smooth on $(0,\infty)$ and has the asymptotic expansion near $u=0$:
$g(u)\sim b_{-1}u^{-1}+\sum_{j=0}^{\infty}\big{(}b_{j}u^{j}-a_{j}\ln u\cdot u^{j}\big{)}.$ (3.18)
Consequently, there exists $\delta>0$, $\varphi,\psi\in C^{\infty}\big{(}[0,\delta)\big{)}$, such that
$g(u)=b_{-1}u^{-1}+\varphi(u)+\psi(u)\ln u,\quad u\in(0,\delta).$ (3.19)
Proof.
From (3.16), it follows that $g(u)\in C^{\infty}((0,\infty))$.
Using (3.14), (3.15), (3.16), and the residue theorem, we obtain for any integer $k\geq 2$,
$\int_{L_{c}}h(s)u^{-s}\,ds-\int_{L_{-k-\frac{1}{2}}}h(s)u^{-s}\,ds=b_{-1}u^{-1}+\sum_{j=0}^{k}\big{(}b_{j}u^{j}-a_{j}\ln u\cdot u^{j}\big{)}.$ (3.20)
Applying (3.16) and (3.17), we obtain the estimate
$\bigg{|}\int_{L_{-k-\frac{1}{2}}}h(s)u^{-s}ds\bigg{|}\leq(2C_{-k+\frac{1}{2},2}+2M_{k})u^{k+\frac{1}{2}}.$ (3.21)
The asymptotic expansion (3.18) follows directly from (3.20) and (3.21). ∎
Lemma 3.16.
Let $v\in C_{c}^{\infty}([0,\infty))$, and let $\hat{v}(s)$ be its Mellin transform:
$\hat{v}(s)=\int_{0}^{\infty}t^{s-1}v(t)dt.$
Then for any integer $m\geq-1$, the function $\hat{v}(s)\Gamma(s+m)$ satisfies the conditions of 3.15.
Proof.
By the standard properties of the Gamma function and the Mellin transform, $\hat{v}(s)\Gamma(s+m)$ has at most double poles at $s=0,-1,-2,\dots$ and a simple pole at $s=1$.
For any integer $k\geq 2$, when ${\mathrm{Re}}(s)>-k$, we have
$\hat{v}(s)=\frac{1}{s(s+1)\cdots(s-1+k)}\int_{0}^{\infty}v^{(k)}(t)t^{s-1+k}dt.$ (3.22)
Furthermore, from Stirling’s formula, for $s\in S_{N,c}$,
$|\Gamma(s+m)|\leq C_{m,N}e^{-\frac{\pi}{4}|\mathrm{Im}(s)|}.$ (3.23)
The bounds (3.16) and (3.17) follow directly from (3.22) and (3.23). ∎
Let $\triangle:=\{({\mathbf{z}},{\mathbf{z}}):{\mathbf{z}}\in U\}$ denotes the diagonal of $U\times U$. The blow up of $U\times U$ is denoted by $\mathrm{Bl}_{\triangle}(U\times U)$, and we have a canonical map $p:\mathrm{Bl}_{\triangle}(U\times U)\rightarrow U\times U$, such that $p^{-1}(\triangle)$ is the exceptional divisor on $\mathrm{Bl}_{\triangle}(U\times U)$.
The following proposition is the main result of this subsection:
Proposition 3.17.
If $C\in S_{\mathrm{reg}}(U\times U)$, then we have
  1. (1)
    $\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C$ has divisorial type singularities along $p^{-1}(\triangle)$ (see B.12). Moreover, any holomorphic derivatives of $\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C$ have divisorial type singularities along $p^{-1}(\triangle)$.
  2. (2)
    $\triangle^{*}\Big{(}\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C\Big{)}:=\lim_{\epsilon\rightarrow 0}\triangle^{*}\Big{(}\int_{\epsilon}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C\Big{)}$
    is a smooth bundle-valued differential form on $U$, where $\triangle^{*}$ is the pull back of bundle-valued differential forms along the diagonal embedding
    $\triangle:U\rightarrow U\times U.$
    Moreover, the pull backs of arbitrary holomorphic derivatives of $\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C$ along $\triangle$ are smooth bundle-valued differential forms.
Proof.
[…]
$\triangle^{*}\left(\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C\right):=\lim_{\epsilon\rightarrow 0}\triangle^{*}\left(\int_{\epsilon}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C\right)$
is a smooth bundle-valued differential form on $U$.
Now, let’s prove $\int_{0}^{\infty}\eta(t)e^{-\frac{\rho^{2}}{2t}}C$ has divisorial type singularities. By Proposition 3.14, we only need to prove that $\tilde{y}_{i}$, $d(\tilde{y}_{i})$, and $\ln(\rho^{2})$ have divisorial type singularities along $p^{-1}(\triangle)$.
Without loss of generality, we assume $U=\mathbb{C}^{n}$. We only need to prove that there’s a neighborhood $W\subset\mathbb{C}^{n}\times\mathbb{C}^{n}$ at $(0,0)$, such that our claims hold in $p^{-1}(W)\subset\mathrm{Bl}_{\triangle}(\mathbb{C}^{n}\times\mathbb{C}^{n})$.
One can identify $\mathrm{Bl}_{\triangle}(\mathbb{C}^{n}\times\mathbb{C}^{n})$ with the closure of the following natural embedding:
$i:\mathbb{C}^{n}\times\mathbb{C}^{n}\setminus\triangle\hookrightarrow\left(\mathbb{C}^{n}\times\mathbb{C}^{n}\right)\times\left(\mathbb{C}^{n}\times\mathbb{C}^{n}\setminus\triangle\right)/\mathbb{C}^{*},\quad({\mathbf{z}},{\mathbf{w}})\mapsto({\mathbf{z}},{\mathbf{w}},{\mathbf{z}},{\mathbf{w}}),$
where the $\mathbb{C}^{*}$-action is given by $\lambda\cdot({\mathbf{z}},{\mathbf{w}})=(\lambda{\mathbf{z}},\lambda{\mathbf{w}})$ for $\lambda\in\mathbb{C}^{*}$.
It is easy to see that the map
$({\mathbf{z}},{\mathbf{w}})\mapsto({\mathbf{w}},\;[z_{1}-w_{1}:\cdots:z_{n}-w_{n}])$
yields an identification of $(\mathbb{C}^{n}\times\mathbb{C}^{n}\setminus\triangle)/\mathbb{C}^{*}$ with $\mathbb{C}^{n}\times\mathbb{P}^{n-1}$. Under this identification, $\mathrm{Bl}_{\triangle}(\mathbb{C}^{n}\times\mathbb{C}^{n})$ has the following description:
$\mathrm{Bl}_{\triangle}(\mathbb{C}^{n}\times\mathbb{C}^{n})$
$\cong\big{\{}({\mathbf{z}},{\mathbf{w}},[\lambda_{1}:\cdots:\lambda_{n}])\in\left(\mathbb{C}^{n}\times\mathbb{C}^{n}\right)\times\mathbb{P}^{n-1}:{\mathbf{z}}-{\mathbf{w}}=k\cdot\boldsymbol{\lambda}\text{ for some }k\in\mathbb{C}\big{\}},$
where $\boldsymbol{\lambda}=(\lambda_{1},\cdots,\lambda_{n})\in\mathbb{C}^{n}.$
Let $\{U_{i}\}_{i=1}^{n}$ be an open cover of $\mathrm{Bl}_{\triangle}(\mathbb{C}^{n}\times\mathbb{C}^{n})$, where each $U_{i}$ corresponds to the subset defined by $\lambda_{i}\neq 0$.
On the chart $U_{1}$, we may use the following local coordinates:
$(z_{1}-w_{1},\lambda_{2},\ldots,\lambda_{n},w_{1},\ldots,w_{n}).$
By the proof of Lemma 2.24, there exists a small neighborhood $W\subset\mathbb{C}^{n}\times\mathbb{C}^{n}$ of $(0,0)$, such that $\frac{\rho^{2}({\mathbf{z}},\bar{\mathbf{z}},{\mathbf{w}},\bar{\mathbf{w}})}{|z_{1}-w_{1}|^{2}}$ is a positive smooth function on $p^{-1}(W)\cap U_{1}$. On the other hand, on $p^{-1}(W)\cap U_{1}$, we have
$\begin{cases}\tilde{y}_{i}&=\frac{1}{z_{1}-w_{1}}\cdot\frac{\bar{\lambda}^{i}|z_{1}-w_{1}|^{2}}{\rho^{2}({\mathbf{z}},\bar{{\mathbf{z}}},{\mathbf{w}},\bar{{\mathbf{w}}})},\\ d(\tilde{y}_{i})&=\frac{d(z_{1}-w_{1})}{(z_{1}-w_{1})^{2}}\cdot\frac{\bar{\lambda}^{i}|z_{1}-w_{1}|^{2}}{\rho^{2}({\mathbf{z}},\bar{{\mathbf{z}}},{\mathbf{w}},\bar{{\mathbf{w}}})}+\frac{1}{z_{1}-w_{1}}\cdot d\Big{(}\frac{\bar{\lambda}^{i}|z_{1}-w_{1}|^{2}}{\rho^{2}({\mathbf{z}},\bar{{\mathbf{z}}},{\mathbf{w}},\bar{{\mathbf{w}}})}\Big{)},\\ \ln(\rho^{2})&=\ln\Big{(}\frac{\rho^{2}}{|z_{1}-w_{1}|^{2}}\Big{)}+\ln(|z_{1}-w_{1}|^{2}).\end{cases}$
Therefore, they have divisorial type singularities on $p^{-1}(W)\cap U_{1}$. Similar results hold on other coordinate charts. This proves our assertion. ∎
Proposition 3.17 reduces the proof of Theorem 2.11 to showing that propagator in the Schwinger space admits a regular expression in a neighborhood of the diagonal. We will prove this in next section.
4 Regularities of Propagator in Schwinger Space
This section establishes the regularity of the propagator in Schwinger space. In Section 4.1, we review the notion of heat kernel expansions. In Section 4.2, we prove 4.3 using Getzler’s rescaling technique.
Let $(M,E,\omega)$ be a triple as in 2.11.
Definition 4.1.
The linear space of regular expressions $S_{\mathrm{reg}}\subset\Omega^{*}\Big{(}(0,\infty);\Gamma\big{(}{\widetilde{E}}\boxtimes{\widetilde{E}}\big{)}\Big{)}$ is generated by differential forms
$w\in\Omega^{*}\Big{(}(0,\infty);\Gamma\big{(}{\widetilde{E}}\boxtimes{\widetilde{E}}\big{)}\Big{)}$,
such that for each local normal frame $(U,{\mathbf{z}},{\mathbf{e}})$, we have $w|_{U\times U}\in S_{\mathrm{reg}}(U\times U)$. Here, we identify $U$ with an open subset of $\mathbb{C}^{n}$, and view $E|_{U}$ as a trivial bundle via the chosen frame.
Definition 4.2.
[…]
It can be seen easily that $\nabla_{\frac{\partial}{\partial z^{i}}},\bar{\partial}\in\mathrm{Op}\big{(}S_{\mathrm{reg}}\big{)}.$
Recall that the propagator in Schwinger space is given by
[…]
The main theorem in this section is that
Theorem 4.3.
We have $e^{\frac{\rho^{2}}{2t}}P_{t}\in S_{\mathrm{reg}}$, where $\rho$ is the distance function induced by the metric $g$.
Fix a local Kähler normal frame $(U,{\mathbf{z}},{\mathbf{e}})$ as described in Appendix A at some point $p\in M$. Let $(U,\mathbf{w},\bf{f})$ denote the same coordinate chart on another copy of $M$.
In this section, without loss of generality, we use this frame to identify $U$ with an open set in $\mathbb{C}^{n}$, equipped with an analytic Kähler metric $g_{i\bar{j}}$, and take $E|_{U}$ to be the trivial holomorphic vector bundle $E=U\times V$, endowed with an analytic Hermitian metric $h_{a\bar{b}}$.
Then under this identification, the heat equation (2.4) takes the more explicit form:
$\partial_{t}H_{t}({\mathbf{z}},\bar{\mathbf{z}},{\mathbf{w}},\bar{\mathbf{w}})$ $=-(\Delta_{\widetilde{E}}\otimes\mathrm{id})H_{t}({\mathbf{z}},\bar{\mathbf{z}},{\mathbf{w}},\bar{\mathbf{w}}),$ (4.1)
$H_{0}({\mathbf{z}},\bar{\mathbf{z}},{\mathbf{w}},\bar{\mathbf{w}})$ $=\delta({\mathbf{z}}-{\mathbf{w}},\bar{{\mathbf{z}}}-\bar{{\mathbf{w}}})\,\omega({\mathbf{z}},{\mathbf{w}})\,d^{n}(\bar{{\mathbf{z}}}-\bar{{\mathbf{w}}}),$ (4.2)
where $\omega({\mathbf{z}},{\mathbf{w}})$ is some holomorphic section of $E\boxtimes E$, $\delta$ denotes the standard delta distribution, and $d^{n}(\bar{{\mathbf{z}}}-\bar{{\mathbf{w}}})=d({\bar{z}}_{1}-{\bar{w}}_{1})\wedge\cdots\wedge d({\bar{z}}_{1}-{\bar{w}}_{1}).$
Since being a regular expression is a local property, it suffices to prove 4.3 on $U\times U$.
4.1 Heat kernel expansion
We aim to prove 4.3 using heat kernel expansions and Getzler’s rescaling technique. In this subsection, we briefly review the relevant properties of heat kernel expansions.
Fix a local Kähler normal frame $(U,{\mathbf{z}},{\mathbf{e}})$ as in Appendix A at some point $p\in M$. Let $(U,\mathbf{w},\bf{f})$ denote the same coordinate chart on another copy of $M$. Let
$\begin{cases}{\tilde{z}}_{i}=z_{i}-w_{i},\\ {\bar{\tilde{z}}}_{i}=\bar{z}_{i}-\bar{w}_{i},\end{cases}$ (4.3)
and let ${\tilde{{\mathbf{z}}}}=({\tilde{z}}_{1},\cdots,{\tilde{z}}_{n})$. Then we will focus on the coordinates
$({\tilde{{\mathbf{z}}}},{\mathbf{w}})$ (4.4)
on $U\times U$.
We have the following heat kernel expansion:
Lemma 4.4.
[…]
  1. (1)
    $u_{i}$ is analytic.
  2. (2)
    For any nonnegative integers $k,l,m$,
    $\bigg{\|}\frac{\partial^{k}}{\partial t^{k}}\Big{(}t^{n}e^{\frac{\rho^{2}}{2t}}H_{t}-\sum_{i=0}^{m}t^{i}u_{i}\Big{)}\bigg{\|}_{C^{l}(K\times K)}\leq C(K,l)t^{m+1-k},$
    where $K\subseteq U$ is a compact subset and $C(K,l)>0$ is a constant. Here
    $\|\cdot\|_{C^{l}(K\times K)}$ denotes the standard $C^{l}$-norm over $K\times K$.
Proof.
The second statement follows directly from [30, Theorem 1.1]. Here we briefly explain the proof of the first statement.
It follows from (4.2) that
$u_{0}({\mathbf{z}},\bar{\mathbf{z}},{\mathbf{w}},\bar{\mathbf{w}})\big{|}_{\triangle}=\omega({\mathbf{z}},{\mathbf{w}})\,d^{n}(\bar{\mathbf{z}}-\bar{\mathbf{w}})|_{\triangle},$ (4.5)
where $\triangle\subset U\times U$ denotes the diagonal.
The heat kernel coefficients $u_{i}$ can be computed recursively by solving ODEs, as described in [4, Chapter 2]. Since both $(M,g)$ and the bundle $E\to M$ are real analytic, and the initial condition is given by (4.5), it follows from standard theory of ODEs that each $u_{i}$ is analytic.
It follows from 4.4 that:
Corollary 4.5.
We have $H_{t}=e^{-\frac{\rho^{2}}{2t}}\Big{(}\big{(}\sum_{i=0}^{n}t^{-n+i}u_{i}\big{)}+b_{0}\Big{)}$ with $t^{-1}b_{0}\in S_{\mathrm{reg}}$.
Proof.
From 4.4, we have when $t\in(0,1)$
$\Big{\|}\frac{\partial^{k}}{\partial t^{k}}\Big{(}e^{\frac{\rho^{2}}{2t}}H_{t}-\sum_{i=0}^{n+k}t^{i-n}u_{i}\Big{)}\Big{\|}_{C^{l}(K\times K)}\leq C_{K,k,l}t.$ (4.6)
This implies that the $k$-th derivatives in the time variable and any spatial derivatives of
$e^{\frac{\rho^{2}}{2t}}H_{t}-\sum_{i=0}^{n+k}t^{i-n}u_{i}$
are well-defined.
Since $\sum_{i=n+1}^{n+k}t^{i-n}u_{i}$ is smooth, the $k$-th time derivatives (for any $k$) and any spatial derivatives of
$b_{0}:=e^{\frac{\rho^{2}}{2t}}H_{t}-\sum_{i=0}^{n}t^{i-n}u_{i}$
also exist.
Finally, by (4.6), we deduce that $t^{-1}b_{0}\in S_{\mathrm{reg}}$. ∎
4.2 Getzler’s rescaling
Our goal in this subsection is to study the regularity properties of the heat kernel $H_{t}$ and the propagator $P_{t}$ using Getzler’s rescaling technique.
We introduce the following (Getzler’s) rescaling $\delta_{\epsilon}$ associated with Kähler normal frame $(U,{\mathbf{z}},{\mathbf{e}})$ and (4.3) as in (3.1):
$\delta_{\epsilon}\big{(}a(\tilde{{\mathbf{z}}},{\bar{\tilde{{\mathbf{z}}}}},t,{\mathbf{w}},\bar{{\mathbf{w}}},d{\bar{\tilde{{\mathbf{z}}}}},d\bar{{\mathbf{w}}},dt)\big{)}=a(\tilde{{\mathbf{z}}},\epsilon{\bar{\tilde{{\mathbf{z}}}}},\epsilon t,{\mathbf{w}},\bar{{\mathbf{w}}},\epsilon d{\bar{\tilde{{\mathbf{z}}}}},d\bar{{\mathbf{w}}},\epsilon dt),$ (4.7)
for any $a\in A\big{(}N(U,r);p\big{)}$.
Note that by 4.4, $\sum_{i=0}^{n}t^{i-n}u_{i}\in A\big{(}N(U,r);p\big{)}.$
To analyze the action of $\delta_{\epsilon}$ on the sum $\sum_{i=0}^{n}t^{i-n}u_{i}$, we consider the following decompositions associated with $(U,{\mathbf{z}},{\mathbf{e}})$ and (4.3).
First, it is clear that each $u_{i}$ admits the decomposition
$u_{i}=\sum_{{\mathbf{j}},{\mathbf{l}}\in\{0,1\}^{n}}u_{i,{\mathbf{j}},{\mathbf{l}}}\,d{\bar{\tilde{{\mathbf{z}}}}}^{{\mathbf{j}}}\wedge d\bar{{\mathbf{w}}}^{{\mathbf{l}}},$
where $u_{i,{\mathbf{j}},{\mathbf{l}}}\in\Gamma(E\boxtimes E)$. Here, for a multi-index ${\mathbf{i}}=(i_{1},\dots,i_{n})\in\{0,1\}^{n}$ such that $i_{j_{1}}=\cdots=i_{j_{\ell}}=1$, we define
$d{\mathbf{z}}^{{\mathbf{i}}}:=dz_{j_{1}}\wedge\cdots\wedge dz_{j_{\ell}}.$
Further decomposing $u_{i,{\mathbf{j}},{\mathbf{l}}}$, we obtain:
$\sum_{i=0}^{n}t^{i-n}u_{i}$ $=\sum_{i=0}^{n}t^{i-n}\Big{(}\sum_{m=0}^{n-i}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=m}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}\,{\bar{\tilde{{\mathbf{z}}}}}^{{\mathbf{k}}}\,d{\bar{\tilde{{\mathbf{z}}}}}^{{\mathbf{j}}}\wedge d\bar{{\mathbf{w}}}^{{\mathbf{l}}}$ (4.8)
$\qquad+\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=n-i+1}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}\,{\bar{\tilde{{\mathbf{z}}}}}^{{\mathbf{k}}}\,d{\bar{\tilde{{\mathbf{z}}}}}^{{\mathbf{j}}}\wedge d\bar{{\mathbf{w}}}^{{\mathbf{l}}}\Big{)},$ (4.9)
with ${\mathbf{j}},{\mathbf{l}}\in\{0,1\}^{n}$ and ${\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{n}$, where
  1. (a)
    if $|{\mathbf{k}}|+|{\mathbf{l}}|\leq n-i$, then $b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}\in\Gamma(E\boxtimes E)$ is holomorphic in the $\tilde{{\mathbf{z}}}$} variable;
  2. (b)
    if $|{\mathbf{k}}|+|{\mathbf{l}}|>n-i$, then $b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}\in\Gamma(E\boxtimes E)$ (not necessarily holomorphic in $\tilde{{\mathbf{z}}}$).
Here for ${\mathbf{j}}=(j_{1},\cdots,j_{n})\in\mathbb{Z}_{\geq 0}^{n}\text{ or }\{0,1\}^{n},|{\mathbf{j}}|:=j_{1}+\cdots+j_{n}$.
We can rearrange the above expansion as follows:
Proposition 4.6.
We have
$\sum_{i=0}^{n}t^{i-n}u_{i}=C_{1}+\sum_{l=-n}^{0}C_{l}$
where
$C_{l}=\sum_{i=0}^{n}t^{i-n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=l+n-i}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{k}}d{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{j}}\wedge d{\bar{\mathbf{w}}}^{\mathbf{l}},l\leq 0$
and
$C_{1}=\sum_{i=0}^{n}t^{i-n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=n-i+1}\sum_{{\mathbf{l}}}b_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}^{i}{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{k}}d{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{j}}\wedge d{\bar{{\mathbf{w}}}}^{\mathbf{l}}.$
To prove 4.3, by 4.5, we need to study $\sum_{l=-n}^{0}C_{l}$ and $C_{1}$. By 3.12, we localize our study at $U\times\{p\}.$ We note that the following lemma, which would be proved in section 4.2.2.
Lemma 4.7.
At $U\times\{p\}$, the limit
$L=\lim_{\epsilon\rightarrow 0}\epsilon\delta_{\epsilon}\left(\partial_{t}+\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)\delta_{\epsilon}^{-1}$ (4.10)
exists.
Then one can see that $C_{l},l\leq 0$ satisfies the following equation:
Proposition 4.8.
Let $q_{t}:=e^{-\frac{|\tilde{{\mathbf{z}}}|^{2}}{2t}}$. Restricted on $U\times\{p\}$,
$L\big{(}q_{t}C_{l}\big{)}=0,\,l\leq 0.$
Here, $L$ denotes the operator defined in (4.10).
Proof.
Since, by the construction above, $\delta_{\epsilon}C_{l}=C_{l}\epsilon^{l}$, and $\delta_{\epsilon}C_{1}=O(\epsilon)$, we have
$\delta_{\epsilon}\Big{(}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}=\sum_{l=-n}^{0}C_{l}\epsilon^{l}+O(\epsilon).$ (4.11)
Following the standard theory of heat kernel expansions (see, e.g., [4, Chapter 2]), we have
$\left(\partial_{t}+\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)\Big{(}e^{-\frac{\rho^{2}}{2t}}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}=e^{-\frac{\rho^{2}}{2t}}\left(\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)u_{n}.$ (4.12)
It follows from (A.1) that, restricted to $U\times\{p\}$,
$\delta_{\epsilon}\left(\frac{\rho^{2}}{2t}\right)=\frac{|\tilde{{\mathbf{z}}}|^{2}}{2t}+O(\epsilon).$ (4.13)
Restricted to $U\times\{p\}$, and using 4.7, together with (4.11)–(4.13), we compute:
$L\bigg{(}q_{t}\Big{(}\sum_{l=-n}^{0}C_{l}\epsilon^{l}\Big{)}\bigg{)}$ $=L\Big{(}\delta_{\epsilon}e^{-\frac{\rho^{2}}{2t}}+O(\epsilon)\Big{)}\bigg{(}\delta_{\epsilon}\Big{(}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}+O(\epsilon)\bigg{)}$
$=L\Big{(}\delta_{\epsilon}e^{-\frac{\rho^{2}}{2t}}\Big{)}\delta_{\epsilon}\Big{(}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}+O(\epsilon)$
$=\left(\epsilon\,\delta_{\epsilon}\left(\partial_{t}+\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)\delta_{\epsilon}^{-1}\right)\delta_{\epsilon}\Big{(}e^{-\frac{\rho^{2}}{2t}}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}+o(1)$
$=\epsilon\,\delta_{\epsilon}\Big{(}e^{-\frac{\rho^{2}}{2t}}\left(\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)u_{n}\Big{)}+o(1)$
$=o(1).$
Examining the order in $\epsilon$, the proposition follows.
The proposition below will be proved in section 4.2.3.
Proposition 4.9.
Let $q_{t}:=e^{-\frac{|\tilde{{\mathbf{z}}}|^{2}}{2t}}$. Suppose $S\in\Omega^{*}\big{(}(0,\infty);\Gamma({\widetilde{E}}\boxtimes{\widetilde{E}}|_{U\times U})\big{)}$ admits an expansion of the form
$S=q_{t}\Big{(}\sum_{i=0}^{n}t^{i-n}v_{i}\Big{)},\quad v_{i}\in\Gamma(\widetilde{E}\boxtimes\widetilde{E}|_{U\times U}).$
Moreover, restricted to $U\times\{p\}$, $S$ is a solution to the equation $LS=0$. Then, if $v_{0}=0$ on $\{p\}\times\{p\}$, we must have $S\equiv 0$ on $U\times\{p\}$. Here, $L$ denotes the operator defined in (4.10).
As a result, we have:
Corollary 4.10.
Restricted on $U\times\{p\}$, we have $C_{l}=0$ if $l<0$.
Proof.
Let $q_{t}:=e^{-\frac{|\tilde{{\mathbf{z}}}|^{2}}{2t}}$. By 4.8,
$L\left(q_{t}C_{l}\right)=0,l\leq 0$
Note that $C_{l}=\sum_{i=0}^{n}t^{i-n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=l+n-i}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{k}}d{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{j}}\wedge d{\bar{\mathbf{w}}}^{\mathbf{l}}$, the coefficient of $t^{-n}$ is
$C_{l}^{0}:=\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=l+n}\sum_{{\mathbf{l}}}b^{0}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{k}}d{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{j}}\wedge d{\bar{\mathbf{w}}}^{\mathbf{l}}.$
However, by (4.2), associated with the frame $(U,{\mathbf{z}},{\mathbf{e}})$,
$u_{0}|_{\{p\}\times\{p\}}$ $=\omega(z,w)d^{n}(\bar{\mathbf{z}}-\bar{\mathbf{w}})|_{\{0\}\times\{0\}}.$ (4.14)
Thus $C_{l}^{0}|_{\{p\}\times\{p\}}=0$ for $l<0$ (since in this case, the multi-index $\mathbf{j}$} in the summation defining $C_{l}^{0}$ satisfies $|{\mathbf{j}}|\leq n+l<n$). Therefore, by 4.9, it follows that $C_{l}=0$ for all $l<0$. ∎
By 4.10 and 4.6, restricted on $U\times\{p\}$,
$\sum_{i=0}^{n}t^{i-n}u_{i}=C_{0}+C_{1}.$ (4.15)
To prove 4.3, we need to deal with the terms $C_{0}$ and $C_{1}$. For this purpose, the following lemma is needed:
Lemma 4.11.
For any multi-index ${\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{n},{\mathbf{j}}\in\{0,1\}^{n}$, we have
  1. (1)
    We have ${\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{k}}d{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{j}}=t^{|{\mathbf{k}}|+|{\mathbf{j}}|}{\mathbf{y}}^{\mathbf{k}}d{\mathbf{y}}^{\mathbf{j}}+t^{|{\mathbf{k}}|+|{\mathbf{j}}|-1}\sum_{j=1}^{n}dt\wedge\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}{\mathbf{y}}^{\mathbf{k}}d{\mathbf{y}}^{\mathbf{j}}.$
  2. (2)
    We have $\sum_{j=1}^{n}\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}\in\mathrm{Op}(S_{{\mathrm{reg}}}).$
Proof.
We first prove (2). Note that $\sum_{i=1}^{n}\iota_{{\bar{\tilde{z}}}_{i}\frac{\partial}{\partial{\bar{z}}_{i}}}$ satisfies the Leibniz rule. It suffices to verify the following identities:
$\sum_{j=1}^{n}\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}dy_{i}=y_{i},\quad\sum_{j=1}^{n}\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}dt=0,$ (4.16)
$\sum_{j=1}^{n}\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}t=\sum_{j=1}^{n}\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}y_{i}=0,$ (4.17)
which are straightforward to verify.
For (1), just note that
$d\bar{\tilde{z}}_{i}=y_{i}\,dt+t\,dy_{i},$
$\sum_{j=1}^{n}\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}\,dy_{i}=y_{i},$
and
$y_{i}\,dt\wedge y_{j}\,dt=0.$
It follows from 4.11 that
Proposition 4.12.
We have
$C_{0}=\Big{(}1+\sum_{j=1}^{n}\frac{dt}{t}\wedge\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}\Big{)}\tilde{C}_{0}=\Big{(}1+\sum_{j=1}^{n}{dt}\wedge y^{j}\iota_{\frac{\partial}{\partial{\bar{z}}_{j}}}\Big{)}\tilde{C}_{0},$
where
$\tilde{C}_{0}=\sum_{i=0}^{n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=n-i}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}{\mathbf{y}}^{\mathbf{k}}d{\mathbf{y}}^{\mathbf{j}}\wedge d{\bar{\mathbf{w}}}^{\mathbf{l}}\in S_{\mathrm{reg}}.$ (4.18)
Proof.
By 4.11 and note that $\iota_{\frac{\partial}{\partial{\bar{z}}^{i}}}{d\bar{\mathbf{w}}}^{\mathbf{l}}=0$, we compute
$\begin{split}C_{0}&=\sum_{i=0}^{n}t^{i-n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=n-i}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{k}}d{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{j}}\wedge d{\bar{\mathbf{w}}}^{\mathbf{l}}\\ &=\sum_{i=0}^{n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=n-i}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}\Big{(}{\mathbf{y}}^{\mathbf{k}}d{\mathbf{y}}^{\mathbf{j}}+\sum_{j=1}^{n}\frac{dt}{t}\wedge\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}{\mathbf{y}}^{\mathbf{k}}d{\mathbf{y}}^{\mathbf{j}}\Big{)}\wedge d{\bar{\mathbf{w}}}^{\mathbf{l}}\\ &=\Big{(}1+\sum_{j=1}^{n}\frac{dt}{t}\wedge\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}\Big{)}\sum_{i=0}^{n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=n-i}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}{\mathbf{y}}^{\mathbf{k}}d{\mathbf{y}}^{\mathbf{j}}\wedge d{\bar{\mathbf{w}}}^{\mathbf{l}}.\end{split}$
As a result,
Proposition 4.13.
There exists $a_{1},a_{2}\in S_{\mathrm{reg}}$, such that restricted on $U\times\{p\}$,
$e^{\frac{\rho^{2}}{2t}}H_{t}$ $={C}_{0}+ta_{1}+dta_{2}$ (4.19)
$=\Big{(}1+\sum_{j=1}^{n}{dt}\wedge y^{j}\iota_{\frac{\partial}{\partial{\bar{z}}_{j}}}\Big{)}\tilde{C}_{0}+ta_{1}+dta_{2}.$
Proof.
First, by 4.5 and (4.15), on $U\times\{p\},$
$H_{t}=e^{-\frac{\rho^{2}}{2t}}\Big{(}\sum_{i=0}^{n}t^{i-n}u_{i}+b_{0}\Big{)}=e^{-\frac{\rho^{2}}{2t}}\left(C_{0}+C_{1}+b_{0}\right).$
Next, proceeding as in 4.12, we have
$C_{1}=\Big{(}1+\sum_{j=1}^{n}\frac{dt}{t}\wedge\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{i}}}\Big{)}\tilde{C}_{1}\quad\text{with\quad$t^{-1}\tilde{C}_{1}\in S_{\mathrm{reg}}.$}$
By (2) in 4.11 and the fact that $\frac{b_{0}}{t}\in{S}_{{\mathrm{reg}}}$ (see 4.5), we can set:
$a_{1}=t^{-1}\tilde{C}_{1}+t^{-1}b_{0}$
and
$a_{2}=\sum_{j=1}^{n}\iota_{{\bar{\tilde{z}}}_{j}\frac{\partial}{\partial{\bar{z}}_{j}}}t^{-1}\tilde{C}_{1}.$
Together with 4.12, the result follows. ∎
Next, we would like to deal with the term $dt\wedge(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id})H_{t}$ in $P_{t}$.
Lemma 4.14.
Let ${\mathcal{M}}=e^{\frac{\rho^{2}}{2t}}(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id})e^{-\frac{\rho^{2}}{2t}}$ and
${{\mathcal{N}}}:=\sum_{i=1}^{n}{y_{i}}\iota_{\frac{\partial}{\partial{\bar{z}}_{i}}}-2\sum_{i=1}^{n}\iota_{\frac{\partial}{\partial{\bar{z}}_{i}}}\frac{\partial}{\partial z_{i}}\in\mathrm{Op}(S_{\mathrm{reg}}).$
Restricted on $U\times\{p\}$, ${{\mathcal{M}}}={{\mathcal{N}}}.$ Thus for any $S\in S_{\mathrm{reg}}$, there exists $S^{\prime}\in S_{\mathrm{reg}}$ such that restricted on $U\times\{p\}$,
${{\mathcal{N}}}S-{\mathcal{M}}S=S^{\prime}.$ (4.20)
Proof.
Since $t\iota_{\frac{\partial}{\partial{\bar{z}}_{i}}}dy_{j}=\delta_{ij}$, we have
$t\,\iota_{\frac{\partial}{\partial{\bar{z}}_{i}}}\in\mathrm{Op}(S_{\mathrm{reg}}).$ (4.21)
Let ${\mathcal{M}}=e^{\frac{\rho^{2}}{2t}}(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id})e^{-\frac{\rho^{2}}{2t}}$, note that $\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id}=-(g^{-1})^{i\bar{j}}\iota_{\frac{\partial}{\partial{\bar{z}}_{j}}}\nabla^{{\widetilde{E}}}_{\frac{\partial}{\partial z_{i}}}$,
${\mathcal{M}}=\sum_{i,j}\Big{(}\frac{1}{2t}\frac{\partial\rho^{2}}{\partial z_{i}}(g^{-1})^{i\bar{j}}\iota_{\frac{\partial}{\partial{\bar{z}}_{j}}}-(g^{-1})^{i\bar{j}}\iota_{\frac{\partial}{\partial{\bar{z}}_{j}}}\nabla^{{\widetilde{E}}}_{\frac{\partial}{\partial z_{i}}}\Big{)}$. (4.22)
It follows from the properties of the Kähler normal frame $(U,{\mathbf{z}},{\mathbf{e}})$ (see Appendix A), together with (4.22) and (4.21), that for any $T\in S_{\mathrm{reg}}$, there exist $T_{i},T_{i}^{\prime},T_{ij}\in S_{\mathrm{reg}}$ such that, when restricted to $U\times\{p\}$, the following identities hold:
$t^{-2}\Big{(}\frac{\partial}{\partial z_{i}}\rho^{2}({\mathbf{z}},\bar{{\mathbf{z}}},{\mathbf{w}},\bar{{\mathbf{w}}})-\frac{\partial}{\partial z_{i}}|\tilde{{\mathbf{z}}}|^{2}\Big{)}T=\Big{(}t^{-2}\frac{\partial}{\partial z_{i}}\rho^{2}({\mathbf{z}},\bar{{\mathbf{z}}},{\mathbf{w}},\bar{{\mathbf{w}}})-t^{-1}y_{i}\Big{)}T=T_{i},$
$t^{-1}\left((g^{-1})^{i\bar{j}}-2\delta^{ij}\right)T=T_{ij},$
and
$t^{-1}\Big{(}\nabla^{\widetilde{E}}_{\frac{\partial}{\partial z_{i}}}-\frac{\partial}{\partial z_{i}}\Big{)}T=t^{-1}\Big{(}h^{-1}\frac{\partial h}{\partial z_{i}}\Big{)}T=T_{i}^{\prime}.$
The last statement then follows.
Proposition 4.15.
There exists $a_{1}^{\prime},a_{2}^{\prime}\in S_{\mathrm{reg}}$ such that restricted on $U\times\{p\}$,
$P_{t}=e^{-\frac{\rho^{2}}{2t}}\bigg{(}\tilde{C}_{0}+2dt\wedge\Big{(}\sum_{i=1}^{n}\iota_{\frac{\partial}{\partial{\bar{z}}_{i}}}\frac{\partial}{\partial z_{i}}\tilde{C}_{0}\Big{)}+ta_{1}^{\prime}+dta_{2}^{\prime}\bigg{)}.$ (4.23)
Proof.
We calculate, restricted on $U\times\{p\}$,
$\quad dt\wedge e^{\frac{\rho^{2}}{2t}}(\bar{\partial}_{\widetilde{E}}^{*}\otimes\mathrm{id})H_{t}$
$=dt\wedge\Big{(}\sum_{i=1}^{n}{y_{i}}\iota_{\frac{\partial}{\partial\bar{z}_{i}}}-2\sum_{i=1}^{n}\iota_{\frac{\partial}{\partial{\bar{z}}_{i}}}\frac{\partial}{\partial z_{i}}\Big{)}\tilde{C}_{0}\mod(dt\wedge{S}_{{\mathrm{reg}}})$
$=C_{0}-\tilde{C}_{0}-2dt\wedge\Big{(}\sum_{i=1}^{n}\iota_{\frac{\partial}{\partial{\bar{z}}_{i}}}\frac{\partial}{\partial z_{i}}\Big{)}\tilde{C}_{0}\mod(dt\wedge{S}_{{\mathrm{reg}}}),$
where the first equality follows from the fact that $(dt)^{2}=0$, along with (4.19) and 4.14; the second equality follows from 4.12.
Note that $P_{t}=-dt\wedge(\bar{\partial}_{{\widetilde{E}}}^{*}\otimes\mathrm{id})H_{t}+H_{t}$, by 4.13 and the computation above, (4.23) follows. ∎
Note that $\tilde{C}_{0}\in S_{\mathrm{reg}}$, to prove 4.3, we still need the following proposition, which will be proved in section 4.2.3.
Proposition 4.16.
Restricted on $U\times\{p\}$,
$\sum_{i=1}^{n}\iota_{\frac{\partial}{\partial{\bar{z}}_{i}}}\frac{\partial}{\partial z_{i}}\tilde{C}_{0}=0.$
4.2.1 Proof of 4.3 and 2.11
By 3.12, to prove 4.3, it suffices to show that for any point $p\in M$, and any Kähler normal frame $(U,{\mathbf{z}},{\mathbf{e}})$ at $p$, there exists $C\in S_{\mathrm{reg}}$ such that on $U\times\{p\}$,
$e^{\frac{\rho^{2}}{2t}}P_{t}=C.$ (4.24)
The identity (4.24) follows directly from 4.15 and 4.16. Therefore, 4.3 is proved.
[…]
4.2.2 Proof of 4.7
The proof of 4.7 follows from the standard argument of Getzler’s rescaling [15, 4]; we include it here for completeness.
Recall that we have fixed a local Kähler normal chart $(U,\mathbf{z},{\mathbf{e}})$ for the bundle $E\to M$, with $\mathbf{z}=(z_{i})_{i=1}^{n}$ centered at some point $p\in M$. Let $(U,\mathbf{w},\mathbf{f})$ denote the same coordinate chart on another copy of $M$.
We use the following coordinates:
$\begin{cases}{\tilde{z}}_{i}=z_{i}-w_{i},\\ {\bar{\tilde{z}}}_{i}=\bar{z}_{i}-\bar{w}_{i}.\end{cases}$ (4.25)
and introduce the rescaling $\delta_{\epsilon}$ associated with these coordinates.
In this subsection, we study the limit over the slice $U\times\{p\}$:
$\lim_{\epsilon\to 0}\epsilon{\delta_{\epsilon}}\left(\partial_{t}+\Delta_{{\widetilde{E}}}\otimes\mathrm{id}\right){\delta_{\epsilon}}^{-1}.$
We adopt the following abbreviations:
  • For vector fields:
    $\partial_{i}:=\frac{\partial}{\partial z_{i}},\text{ and }\partial_{\bar{j}}:=\frac{\partial}{\partial{\bar{z}}_{j}}.$
  • For any connection $\nabla$ on a complex vector bundle $H\to M$:
    $\nabla_{i}:=\nabla_{\partial_{i}},\quad\nabla_{{\bar{j}}}:=\nabla_{\partial_{{\bar{j}}}}.$
  • If $F$ is the curvature of $\nabla$, then:
    $F_{i{\bar{j}}}:=F\left(\partial_{i},\partial_{{\bar{j}}}\right)\in\mathrm{End}(H).$
First, we recall the local expression for the Laplacian:
$\Delta_{{\widetilde{E}}}=-\sum_{i,j}\Big{(}g^{i{\bar{j}}}\nabla^{{\widetilde{E}}}_{\partial_{i}}\nabla^{{\widetilde{E}}}_{\partial_{{\bar{j}}}}+g^{i{\bar{j}}}\nabla^{{\widetilde{E}}}_{\nabla^{TM}_{\partial_{i}}\partial_{{\bar{j}}}}-\sum_{l}d{\bar{z}}_{i}\,\iota_{\partial_{\bar{l}}}F^{E\otimes K}_{j\bar{i}}g^{j\bar{l}}\Big{)},$
where $K$ denotes the canonical bundle, and $F^{E\otimes K}$ is the curvature of $E\otimes K$, induced by the connections $\nabla^{TM}$ and $\nabla^{E}$.
The following identities from Kähler geometry will be useful:
$\nabla^{TM}_{i}\partial_{{\bar{j}}}$ $=0,\quad\nabla^{TM}_{\bar{j}}\partial_{i}=0,$
$\nabla^{TM}_{i}\partial_{j}$ $=\sum_{k}w_{ij}^{k}\partial_{k},\quad\nabla^{TM}_{\bar{i}}\partial_{{\bar{j}}}=\sum_{k}w_{\bar{i}{\bar{j}}}^{\bar{k}}\partial_{\bar{k}},$
$\nabla^{E}_{i}e_{a}$ $=\sum_{b}\Gamma_{ia}^{b}e_{b},\quad\nabla_{{\bar{j}}}^{E}e_{a}=0,$
where
$w_{ij}^{k}=\sum_{l}g^{k\bar{l}}\partial_{i}g_{j\bar{l}},\quad w_{\bar{i}{\bar{j}}}^{\bar{k}}=\sum_{l}g^{l\bar{k}}\partial_{\bar{i}}g_{l\bar{j}},\quad\Gamma_{ia}^{b}=\sum_{c}h^{b\bar{c}}\partial_{i}h_{a\bar{c}}.$ (4.26)
Note that $\nabla^{E}_{{\bar{j}}}=\partial_{{\bar{j}}}$ and $\nabla^{TM}_{i}d{\bar{z}}_{j}=0$. Then we obtain:
$\nabla^{{\widetilde{E}}}_{\bar{j}}$ $=\partial_{{\bar{j}}}+\sum_{i,k}w_{\bar{i}\bar{j}}^{\bar{k}}d{\bar{z}}_{i}\wedge\iota_{\partial_{\bar{k}}}$
$=\partial_{{\bar{j}}}+\sum_{i,k}w_{\bar{i}\bar{j}}^{\bar{k}}(d{\bar{\tilde{z}}}_{i}+d\bar{w}_{i})\wedge\iota_{\partial_{\bar{k}}},$
$\nabla^{{\widetilde{E}}}_{i}$ $=\partial_{i}+\Gamma_{i},$
$\sum_{i,j,l}d{\bar{z}}_{i}\,\iota_{\partial_{\bar{l}}}F_{j\bar{i}}^{E\otimes K}g^{j\bar{l}}$ $=\sum_{i,j,l}(d{\bar{\tilde{z}}}_{i}+d{\bar{w}}_{i})\,\iota_{\partial_{\bar{l}}}F_{j\bar{i}}^{E\otimes K}g^{j\bar{l}},$
where $\Gamma_{i}:=\sum_{a,b}\Gamma_{ia}^{b}e^{a}\otimes e_{b}\in\mathrm{End}(E)$, and $\{e^{a}\}$ is the dual frame of $\{e_{a}\}$.
Note that on $U\times\{p\}$, where ${\tilde{z}}=z$, ${\bar{\tilde{z}}}={\bar{z}}$, we have:
$\Gamma_{ia}^{b}({\mathbf{z}},\bar{\mathbf{z}})=\Gamma_{ia}^{b}({\tilde{{\mathbf{z}}}},\bar{\tilde{{\mathbf{z}}}})=\sum_{j}f_{ija}^{b}({\tilde{{\mathbf{z}}}},\bar{\tilde{{\mathbf{z}}}}){\bar{\tilde{z}}}_{j},$
for some analytic $f_{ija}^{b}$. As a result,
$\delta_{\epsilon}\Gamma_{ia}^{b}=O(\epsilon).$
Thus, at $U\times\{p\}$,
$\epsilon{\delta_{\epsilon}}\nabla_{{\bar{j}}}^{{\widetilde{E}}}{\delta_{\epsilon}}^{-1}$ $=\bar{\partial}_{\bar{j}}+\sum_{i,k}w_{\bar{i}\bar{j}}^{\bar{k}}(\tilde{\mathbf{z}},0)d{\bar{w}}_{i}\iota_{\bar{\partial}_{\bar{k}}}+O(\epsilon),$
${\delta_{\epsilon}}\nabla^{{\widetilde{E}}}_{i}{\delta_{\epsilon}}^{-1}$ $=\partial_{i}+{\delta_{\epsilon}}\Gamma_{i}=\partial_{i}+O(\epsilon),$
${\delta_{\epsilon}}g^{i{\bar{j}}}$ $=2\delta^{ij}+O(\epsilon),$
$\epsilon{\delta_{\epsilon}}\Big{(}\sum_{i,j,l}(d{\bar{\tilde{z}}}_{i}+d{\bar{w}}_{i})\iota_{\bar{\partial}_{\bar{l}}}F_{j\bar{i}}^{E\otimes K}g^{j\bar{l}}\Big{)}$ $=\sum_{i,j}d{\bar{w}}_{i}\iota_{\bar{\partial}_{{\bar{j}}}}F_{j\bar{i}}^{E\otimes K}+O(\epsilon).$
Here, for an analytic function $f$ on $C^{\infty}(\mathbb{C}^{n})$, $f({\mathbf{z}},0)$ denotes its holomorphic component.
Therefore, at $U\times\{p\}$,
$\epsilon{\delta_{\epsilon}}\Delta_{{\widetilde{E}}}{\delta_{\epsilon}}^{-1}=-2\sum_{i}\partial_{i}\Big{(}\bar{\partial}_{\bar{i}}+\sum_{j,k}w_{\bar{j}\bar{i}}^{\bar{k}}({{\mathbf{z}}},0)d{\bar{w}}_{j}\iota_{\bar{\partial}_{\bar{k}}}\Big{)}+2\sum_{i,j}d{\bar{w}}_{i}\iota_{\bar{\partial}_{\bar{j}}}F_{j\bar{i}}^{E\otimes K}({{\mathbf{z}}},0)+O(\epsilon).$ (4.27)
We define
$\Delta^{{\mathrm{md}}}:=-2\sum_{i}\partial_{i}\Big{(}\bar{\partial}_{\bar{i}}+\sum_{j,k}w_{\bar{j}\bar{i}}^{\bar{k}}({{\mathbf{z}}},0)d{\bar{w}}_{j}\iota_{\bar{\partial}_{\bar{k}}}\Big{)}+2\sum_{i,j}d{\bar{w}}_{i}\iota_{\bar{\partial}_{\bar{j}}}F_{j\bar{i}}^{E\otimes K}({{\mathbf{z}}},0).$
Then the operator $L$ appearing in 4.7 is given by
$L=\partial_{t}+\Delta^{{\mathrm{md}}}\otimes\mathrm{id}.$
4.2.3 Proof of 4.9 and 4.16
Let $\nabla^{\mathrm{md}}$ be the operator given by
$\nabla^{\mathrm{md}}_{\bar{j}}=\partial_{\bar{j}}+\sum_{l,k}w_{\bar{l}\bar{j}}^{\bar{k}}(\tilde{\mathbf{z}},0)d{\bar{w}}_{l}\iota_{\partial_{\bar{k}}},\quad\text{and}\quad\nabla^{{\mathrm{md}}}_{j}=\partial_{j}.$
Then the Laplacian $\Delta^{{\mathrm{md}}}$ is the connection Laplacian associated with $\nabla^{{\mathrm{md}}}$:
$\Delta^{{\mathrm{md}}}=-2\sum_{i}\nabla^{{\mathrm{md}}}_{i}\nabla^{{\mathrm{md}}}_{\bar{i}}+2\sum_{i,j}d\bar{w}_{i}\iota_{\bar{\partial}_{\bar{j}}}F_{j\bar{i}}^{E\otimes K}({\mathbf{z}},0).$
Suppose $)S\in\Omega^{*}((0,\infty);\Gamma({\widetilde{E}}\times{\widetilde{E}}$)) satisfies the condition in 4.9. That is, it solves the heat equation at $U\times\{p\}$:
$\big{(}\partial_{t}+\Delta^{\mathrm{md}}\big{)}S=0,$ (4.28)
and admits the following asymptotic expansion at $U\times\{p\}$:
[…]
We compute, at $U\times\{p\}$,
$\Big{(}\partial_{t}+\Delta^{{\mathrm{md}}}\Big{)}S=t^{-n}e^{-\frac{|\tilde{{\mathbf{z}}}|^{2}}{2t}}\bigg{(}\partial_{t}+\Delta^{{\mathrm{md}}}+\sum_{i}\Big{(}\frac{\bar{\tilde{z}}_{i}}{t}\nabla^{{\mathrm{md}}}_{\bar{i}}+\frac{\tilde{z}_{i}}{t}\nabla^{{\mathrm{md}}}_{i}\Big{)}\bigg{)}\sum_{k=0}^{n}t^{k}v_{k}.$ (4.30)
[…]
$\Big{(}\partial_{t}+\Delta^{{\mathrm{md}}}\Big{)}S=t^{-n}e^{-\frac{|\tilde{{\mathbf{z}}}|^{2}}{2t}}\left(\partial_{t}+\Delta^{{\mathrm{md}}}+\frac{\nabla_{r\nabla r}}{t}\right)\sum_{k=0}^{n}t^{k}v_{k}.$ (4.31)
Hence, we obtain the recursive system:
$\begin{cases}\nabla^{{\mathrm{md}}}_{r\nabla r}v_{0}=0,\\ \nabla^{{\mathrm{md}}}_{r\nabla r}v_{k+1}+(k+1)v_{k+1}+\Delta^{{\mathrm{md}}}v_{k}=0,\quad k\geq 0.\end{cases}$ (4.32)
Proof of 4.9.
We focus on the analysis at $U\times\{p\}$.
It follows from standard ODE theory that if $v_{0}=0$ at ${\tilde{{\mathbf{z}}}}=0$ and satisfies (4.32), then $v_{0}\equiv 0$.
Next, since $v_{0}$ already vanishes, the second equation in (4.32) yields
$\nabla^{\mathrm{md}}_{\nabla r}(rv_{1})=0.$
By the same reasoning, since $rv_{1}=0$ at $\tilde{{\mathbf{z}}}=0$, we have $rv_{1}\equiv 0$, and hence $v_{1}\equiv 0$.
Proceeding inductively, suppose $v_{k}\equiv 0$. Then from (4.32), we obtain
$\nabla^{\mathrm{md}}_{\nabla r}(r^{k+1}v_{k+1})=0.$
Since $r^{k+1}v_{k+1}$ vanishes at ${\tilde{{\mathbf{z}}}}=0$, it follows that $v_{k+1}\equiv 0$. Therefore, $v_{k}\equiv 0$ for all $k\geq 0$, completing the proof.
Lemma 4.17.
At $U\times\{p\}$, the following commutation relations hold:
  1. (a)
    $\left[\sum_{l}\iota_{\partial_{\bar{l}}}\partial_{l},\nabla^{\mathrm{md}}_{{j}}\right]=0.$
  2. (b)
    $\left[\sum_{l}\iota_{\partial_{\bar{l}}}\partial_{l},\nabla^{\mathrm{md}}_{\bar{j}}\right]=0.$
  3. (c)
    $\left[\sum_{l}\iota_{\partial_{\bar{l}}}\partial_{l},\Delta^{\mathrm{md}}\right]=0.$
Proof.
We again work at $U\times\{p\}$.
Part (a) is straightforward.
To prove (b), observe that the curvature term satisfies
$R_{\bar{i}l\bar{j}}^{\bar{k}}=\partial_{l}w_{\bar{i}\bar{j}}^{\bar{k}},\quad\text{and}\quad R_{\bar{i}l\bar{j}}^{\bar{k}}=R_{\bar{i}k\bar{j}}^{\bar{l}}.$
Then
$\quad\Big{[}\sum_{l}\iota_{\partial_{\bar{l}}}\partial_{l},\sum_{i,k}w_{\bar{i}\bar{j}}^{\bar{k}}({\tilde{{\mathbf{z}}}},0)d\bar{w}_{i}\iota_{{\partial}_{\bar{k}}}\Big{]}=-\sum_{i,k,l}\partial_{l}w_{\bar{i}\bar{j}}^{\bar{k}}({\tilde{{\mathbf{z}}}},0)d\bar{w}_{i}\iota_{\partial_{\bar{l}}}\iota_{{\partial}_{\bar{k}}}$ (4.33)
$=-\sum_{i,k,l}R_{\bar{i}l\bar{j}}^{\bar{k}}({\tilde{{\mathbf{z}}}},0)d\bar{w}_{i}\iota_{\partial_{\bar{l}}}\iota_{\partial_{\bar{k}}}=-\frac{1}{2}\sum_{i,k,l}R_{\bar{i}l\bar{j}}^{\bar{k}}({\tilde{{\mathbf{z}}}},0)d\bar{w}_{i}[\iota_{\partial_{\bar{l}}},\iota_{\partial_{\bar{k}}}]=0.$
Since
$\nabla^{\mathrm{md}}_{\bar{j}}=\bar{\partial}_{\bar{j}}+\sum_{i,k}w_{\bar{i}\bar{j}}^{\bar{k}}({\tilde{{\mathbf{z}}}},0)d\bar{w}_{i}\iota_{{\partial}_{\bar{k}}},$
equation (4.33) implies part (b).
For (c), recall that
$\sum_{l}\iota_{\partial_{\bar{l}}}\partial_{l}=-\bar{\partial}^{{\mathrm{md}},*},\quad\text{and}\quad\Delta^{\mathrm{md}}=[\bar{\partial}^{{\mathrm{md}}},\bar{\partial}^{{\mathrm{md}},*}].$
Thus,
$\bar{\partial}^{{\mathrm{md}},*},\Delta^{\mathrm{md}}]=0,$
which proves (c). ∎
Consider
$\tilde{v}_{k}:=\sum_{l}\iota_{\partial_{\bar{l}}}\partial_{l}v_{k},\quad k\geq 0.$
By 4.17, the $\tilde{v}_{k}$ satisfy:
$\begin{cases}\nabla^{\mathrm{md}}_{r\nabla r}\tilde{v}_{0}+\tilde{v}_{0}=0,\\ \nabla^{\mathrm{md}}_{r\nabla r}\tilde{v}_{k+1}+(k+2)\tilde{v}_{k+1}+\Delta^{\mathrm{md}}\tilde{v}_{k}=0,\quad k\geq 0,\end{cases}$ (4.34)
[…]
Proof of 4.16.
As in the proof of 4.9, equations (4.34) implies that $\tilde{v}_{k}=0$ for all $k\geq 0$ at $U\times\{p\}$. ∎
Appendix A Existence of Canonical Local Coordinates
We introduce suitable coordinate systems to simplify our discussion, specifically Kähler normal coordinates on the manifold and holomorphic local frames on the vector bundle.
It is standard that
Proposition A.1.
Let $M$ be an $n$-dimensional real analytic Kähler manifold, and let $p\in M$. Then there exists a holomorphic coordinate system $\mathbf{z}=(z_{1},\ldots,z_{n})$ in a neighborhood $U$ of $p$ such that $\mathbf{z}(p)=0$ and
$\left(\frac{\partial^{k}g_{i\bar{j}}}{\partial z_{i_{1}}\cdots\partial z_{i_{k}}}\right)(p)=0,\quad\text{for }k\geq 1,$
and $g_{i\bar{j}}(p)={\frac{1}{2}}\delta_{ij}$, where $g$ denotes the Riemannian metric on $M$ and $g_{i\bar{j}}:=g\left(\frac{\partial}{\partial z_{i}},\frac{\partial}{\partial\bar{z}_{j}}\right)$. The pair $(U,\mathbf{z})$ is called the Kähler normal chart at $p$, and $\mathbf{z}$ is called a Kähler coordinate system at $p$, which is unique up to unitary transformations.
Proof.
The proof can be found in [22, 32]. ∎
We have the following corollary for Kähler normal coordinates:
Corollary A.2.
Let $M$ be an $n$-dimensional real analytic Kähler manifold. Let $p\in M$, and let $(U,\mathbf{z})$ be a Kähler normal chart at $p$. Consider another copy of $M$, and let $(U,\mathbf{w})$ denote the corresponding Kähler normal chart at $p$ on this second copy.
Let $\rho^{2}:M\times M\to\mathbb{R}$ denote the square of the distance function. With respect to the charts above, we write it as $\rho^{2}(\mathbf{z},\bar{\mathbf{z}},\mathbf{w},\bar{\mathbf{w}})$. Then:
  1. (1)
    On $U\times\{p\}$, we have
    $\rho^{2}(\mathbf{z},\bar{\mathbf{z}},0,0)=\sum_{i=1}^{n}z_{i}\bar{z}_{i}+\sum_{i,j,k,l=1}^{n}z_{i}z_{j}\bar{z}_{k}\bar{z}_{l}\,f_{ij\bar{k}\bar{l}}(\mathbf{z},\bar{\mathbf{z}}),$ (A.1)
    where $f_{ij\bar{k}\bar{l}}$ are real analytic functions on $U$.
  2. (2)
    In general, we have
    $\rho^{2}(\mathbf{z},\bar{\mathbf{z}},\mathbf{w},\bar{\mathbf{w}})=\sum_{i=1}^{n}(z_{i}-w_{i})(\bar{z}_{i}-\bar{w}_{i})+\sum_{i,j=1}^{n}(z_{i}-w_{i})(\bar{z}_{j}-\bar{w}_{j})\,f_{i\bar{j}}(\mathbf{z},\bar{\mathbf{z}},\mathbf{w},\bar{\mathbf{w}}),$ (A.2)
    where $f_{i\bar{j}}$ are real analytic functions on $U\times U$ satisfying $f_{i\bar{j}}(0,0,0,0)=0$.
Proof.
We recall the Eikonal equation on Kähler manifolds (see [29, Example 3.2]):
$\sum_{i,j=1}^{n}\partial_{i}\big{(}\rho^{2}({\mathbf{z}},\bar{\mathbf{z}},0,0)\big{)}g^{i\bar{j}}({\mathbf{z}},\bar{\mathbf{z}})\partial_{\bar{j}}\big{(}\rho^{2}({\mathbf{z}},\bar{\mathbf{z}},0,0)\big{)}=2\rho^{2}({\mathbf{z}},\bar{\mathbf{z}},0,0).$ (A.3)
[…]
Proposition A.3.
Let $M$ be an $n$-dimensional real analytic complex manifold, and let $E\to M$ be a real analytic Hermitian holomorphic vector bundle. Let $p\in M$, and let $(U,\mathbf{z})$ be a holomorphic chart centered at $p$, i.e., $\mathbf{z}(p)=0$, such that $E|_{U}\to U$ is trivial.
Then there exists a local holomorphic frame $\mathbf{e}=(e_{a})_{a=1}^{\mathrm{rank}(E)}$ of $E$ over $U$ such that
$\left(\frac{\partial^{k}h_{a\bar{b}}}{\partial z_{i_{1}}\cdots\partial z_{i_{k}}}\right)(p)=0\quad\text{for all }k\geq 1,$
and
$h_{a\bar{b}}(p)=\delta_{ab},$
where $h_{a\bar{b}}:=h(e_{a},e_{b})$.
We refer to $(U,\mathbf{z},\mathbf{e})$ as a holomorphic normal frame at $p$. For each $(U,{\mathbf{z}})$, the frame $\mathbf{e}$} is unique up to unitary transformations. Moreover, if $(U,\mathbf{z})$ is a Kähler normal chart at $p$, then we say that $(U,\mathbf{z},\mathbf{e})$ is a Kähler normal frame at $p$.
Proof.
First, for each holomorphic coordinate system $\mathbf{z}$} such that ${\mathbf{z}}(p)=0$, choose an arbitrary holomorphic frame ${\mathbf{e}}^{\prime}=(e_{a}^{\prime})$ such that $h_{a\bar{b}}(p)=\delta_{ab}$. To establish the proposition, we aim to construct a holomorphic matrix $G_{a}^{b}(\mathbf{z})$ such that
$G_{a}^{b}(\mathbf{z})\,h_{b\bar{c}}(\mathbf{z},\bar{\mathbf{z}})\,\overline{G_{d}^{c}(\mathbf{z})}=\delta_{ad}+\sum_{\mathbf{j}\in\mathbb{Z}_{>0}^{n}}\bar{\mathbf{z}}^{\mathbf{j}}f_{ad,\mathbf{j}}(\mathbf{z},\bar{\mathbf{z}})$ (A.5)
for some smooth $f_{ab,{\mathbf{j}}}$, where for any $\mathbf{j}=(j_{1},\ldots,j_{n})\in\mathbb{Z}_{>0}^{n}$, we set $\bar{\mathbf{z}}^{\mathbf{j}}:=\bar{z}_{1}^{j_{1}}\cdots\bar{z}_{n}^{j_{n}}$. With this matrix, setting $e_{a}=G_{a}^{b}e^{\prime}_{b}$ and ${\mathbf{e}}=(e_{a})$, it follows from (A.5) that $({\mathbf{z}},{\mathbf{e}})$ forms a holomorphic normal frame.
To determine $G_{a}^{b}$, we analytically continue (A.5) to a neighborhood of $(p,\bar{p})\in M\times\bar{M}$ and evaluate it at $(\mathbf{z},0)$, obtaining
$G_{a}^{b}(\mathbf{z})\,h_{b\bar{c}}(\mathbf{z},0)\,\overline{G_{d}^{c}(0)}=\delta_{ad}.$
By setting
$G_{a}^{b}(\mathbf{z})=\left(h^{-1}(\mathbf{z},0)\right)^{b\bar{c}}\delta_{ca},$
we verify that $G_{a}^{b}(\mathbf{z})$ satisfies equation (A.5). ∎
Appendix B Cauchy Principal Value
In this appendix, we will introduce the Cauchy principal value for the reader’s convenience. We will mainly follow [18], with extra attention to logarithmic terms.
Let’s first consider the local case.
Local case
For any ${\mathbf{z}}=(z_{1},\cdots,z_{m+n})\in\mathbb{C}^{m+n}$, ${\mathbf{i}}=(i_{1},\cdots,i_{m})\in\mathbb{Z}^{m}$, we set
${\mathbf{z}}^{\mathbf{i}}=z_{1}^{i_{1}}\cdots z_{m}^{i_{m}}$
and
$\ln({\mathbf{z}},{\mathbf{i}})=\big{(}\ln(|z_{1}|^{2})\big{)}^{i_{1}}\cdots\big{(}\ln(|z_{m}|^{2})\big{)}^{i_{m}}$
Definition B.1.
Given a differential form $\alpha\in\Omega^{*,*}\left(\left(\mathbb{C}^{*}\right)^{m}\times\mathbb{C}^{n}\right)$. We say $\alpha$ is a differential form on $\mathbb{C}^{m+n}$ with divisorial type singularities of $({\mathbf{i}},{\mathbf{k}})$-type along the principal divisor: $f=z_{1}z_{2}\cdots z_{m}$, if there exist $\beta\in\Omega^{*,*}\left(\mathbb{C}^{m+n}\right)$, ${\mathbf{i}}\in\mathbb{Z}_{\geq 0}^{m}$ and $\bf\in\mathbb{Z}_{\geq 0}^{m}$, such that
$\alpha={\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\beta.$
If ${\mathbf{k}}=0$, we say $\alpha$ is logarithm-free.
We will prove in this appendix that
Theorem B.2.
Given the following data:
  1. (a)
    A compactly supported $(m+n,m+n)$ differential form $\alpha$ on $\mathbb{C}^{m+n}$ with divisorial type singularities of $({\mathbf{i}},{\mathbf{k}})$ along $z_{1}\cdots z_{m}$.
  2. (b)
    A holomorphic function ${\mathbf{z}}^{{\mathbf{j}}}$, where ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$.
  3. (c)
    A nowhere vanishing smooth function $f$ on a neighborhood of support of $\alpha$.
Then the following limit exists:
$\lim_{\delta\to 0}\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}\alpha.$ (B.1)
Moreover, the limit is independent of the choice of $\mathbf{j}$} and smooth function $f$.
Definition B.3.
With the same assumptions as Theorem B.2, we call the limit
$\lim_{\delta\to 0}\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}\alpha$
the integral of $\alpha$ over $\mathbb{C}^{m+n}$ in the sense of Cauchy principal value. We will use
[…]
to denote it.
By Theorem B.2 $(c)$, we see that the Cauchy principal value only depends on the support of the divisor, and not on the multiplicity or the choice of the holomorphic function ${\mathbf{z}}^{\mathbf{j}}$ representing it. This feature allows us to define the Cauchy principal value on general complex manifolds.
Before we get into the global case, let us prove Theorem B.2.
Let $B:=\{{\mathbf{z}}\in\mathbb{C}^{m+n}:\sup_{1\leq i\leq m+n}|z_{i}|\leq 1\}.$
For each ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$, let $S^{\mathbf{j}}_{\delta}:=\big{\{}{\mathbf{z}}\in B:|{\mathbf{z}}^{\mathbf{j}}|=\delta\big{\}}$ and $B_{\delta}^{\mathbf{j}}:=\big{\{}{\mathbf{z}}\in B:|{\mathbf{z}}^{\mathbf{j}}|\in(\delta,1)\big{\}}$.
For each $l\in\{1,2,\cdots,m+n\}$, set ${\mathbf{z}}_{{\hat{i}}}=(z_{1},\cdots,z_{l-1},z_{l+1},\cdots,z_{m+n})\in\mathbb{C}^{m+n-1}$ and for each ${l}\in\{1,\cdots,m\}$, ${\mathbf{j}}_{\hat{l}}:=(j_{1},\cdots,j_{l-1},j_{l+1},\cdots,j_{m}).$
Similar to [18, Lemma 6.4], we have
Lemma B.4.
Let $\phi\in L^{\infty}(\mathbb{C}^{m+n})$, then for any ${\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m},{\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$, any $i\in\{1,2,\cdots,m\}$,
$\lim_{\delta\to 0}\int_{S^{\mathbf{j}}_{\delta}}\phi({\mathbf{z}},\bar{\mathbf{z}})\ln({\mathbf{z}},{\mathbf{k}})$ $dz_{i}\wedge d{\mathbf{z}}_{{\hat{i}}}\wedge d\bar{\mathbf{z}}_{{\hat{i}}}=0,$ (B.2)
$\lim_{\delta\to 0}\int_{S^{\mathbf{j}}_{\delta}}\phi({\mathbf{z}},\bar{\mathbf{z}})\ln({\mathbf{z}},{\mathbf{k}})$ $d\bar{z}_{i}\wedge d{\mathbf{z}}_{{\hat{i}}}\wedge d\bar{\mathbf{z}}_{{\hat{i}}}=0.$
Here $d{\mathbf{z}}_{{\hat{i}}}:=dz_{1}\wedge\cdots\wedge dz_{i-1}\wedge dz_{i+1}\cdots\wedge dz_{m+n}$ and $d\bar{{\mathbf{z}}}_{{\hat{i}}}:=\overline{d{\mathbf{z}}_{{\hat{i}}}}$.
Proof.
Let $\mathbf{k}=(k_{1},\dots,k_{m})$ and $\mathbf{j}=(j_{1},\dots,j_{m})$, and denote $|\mathbf{j}|:=j_{1}+\dots+j_{m}$.
Without loss of generality, we may assume that $i=1$. Recall that by the definition of $S_{\delta}^{\mathbf{j}}$, we have $S_{\delta}^{\mathbf{j}}\subset B$. Consider the following parametrization of $S_{\delta}^{\mathbf{k}}$:
$(\theta,{\mathbf{z}}_{\hat{1}})\mapsto\Bigg{(}\bigg{(}\frac{\delta}{\prod_{s=2}^{m}\rho_{s}^{j_{s}}}\bigg{)}^{1/j_{1}}e^{i\theta},{\mathbf{z}}_{\hat{1}}\Bigg{)}$,
where $z_{s}=\rho_{s}e^{i\theta_{s}}$. With this parametrization, one verifies that the integrals appearing in (B.2) are bounded, up to a constant, by
$\delta^{1/j_{1}}\left(\ln\delta\right)^{|\mathbf{k}|}\cdot\int_{E^{\mathbf{j}}_{\delta}}\prod_{s=2}^{m}\rho_{s}^{1-j_{s}/j_{1}}\,d\rho_{s},$ (B.3)
where for each $\mathbf{r}=(r_{1},\cdots,r_{m})\in\mathbb{R}_{>0}^{m},$
$E^{\mathbf{r}}_{\delta}:=\Big{\{}(\rho_{2},\dots,\rho_{m})\in\mathbb{R}_{\geq 0}^{m}:\delta\leq\prod_{s=2}^{m}\rho_{s}^{r_{s}}\leq 1,\quad\sup_{s\in\{2,\cdots,m\}}\rho_{s}\leq 1\Big{\}}.$
We now prove by induction that for any $r\geq 0$ and $\mathbf{r}=(r_{1},\dots,r_{m})\in\mathbb{R}_{>0}^{m}$, the following limit holds:
$\lim_{\delta\to 0}\delta^{1/r_{1}}(\ln\delta)^{r}\cdot\int_{E^{\mathbf{r}}_{\delta}}\prod_{s=2}^{m}\rho_{s}^{1-r_{s}/r_{1}}\,d\rho_{s}=0.$ (B.4)
The base case $m=2$ follows from a direct computation.
Let
$E^{\prime}_{\delta}:=\Big{\{}(\rho_{3},\dots,\rho_{m})\in\mathbb{R}_{\geq 0}^{m}:\delta\leq\prod_{s=3}^{m}\rho_{s}^{r_{s}}\leq 1,\quad\sup_{s\in\{3,\cdots,m\}}\rho_{s}\leq 1\Big{\}}.$
We distinguish two cases:
  1. (a)
    Case $r_{2}\leq 2r_{1}$:
    By Fubini’s theorem,
    $\delta^{1/r_{1}}(\ln\delta)^{r}\cdot\int_{E^{\mathbf{r}}_{\delta}}\prod_{s=2}^{m}\rho_{s}^{1-r_{s}/r_{1}}\,d\rho_{s}$
    $\leq\delta^{1/r_{1}}(\ln\delta)^{r}\cdot\int_{E^{\prime}_{\delta}}\prod_{s=3}^{m}\rho_{s}^{1-r_{s}/r_{1}}\,d\rho_{s}\int_{\delta^{1/r_{2}}}^{1}\rho_{2}^{1-r_{2}/r_{1}}\,d\rho_{2}$
    $\leq C\delta^{1/r_{1}}(\ln\delta)^{r+1}\cdot\int_{E^{\prime}_{\delta}}\prod_{s=3}^{m}\rho_{s}^{1-r_{s}/r_{1}}\,d\rho_{s}.$
    By the induction hypothesis, the right-hand side tends to zero as $\delta\to 0$, hence the limit in (B.4) holds in this case.
  2. (b)
    Case $r_{2}>2r_{1}$:
    Again applying Fubini’s theorem,
    $\delta^{1/r_{1}}(\ln\delta)^{r}\cdot\int_{E^{\mathrm{r}}_{\delta}}\prod_{s=2}^{m}\rho_{s}^{1-r_{s}/r_{1}}\,d\rho_{s}$
    $\leq\delta^{1/r_{1}}(\ln\delta)^{r}\cdot\int_{E^{\prime}_{\delta}}\prod_{s=3}^{m}\rho_{s}^{1-r_{s}/r_{1}}\,d\rho_{s}\int_{\left(\delta/\prod_{s=3}^{m}\rho_{s}^{r_{s}}\right)^{1/r_{2}}}^{+\infty}\rho_{2}^{1-r_{2}/r_{1}}\,d\rho_{2}$
    $\leq C_{1}\delta^{2/r_{2}}(\ln\delta)^{r}\cdot\int_{E^{\prime}_{\delta}}\prod_{s=3}^{m}\rho_{s}^{1-2r_{s}/r_{2}}\,d\rho_{s}$
    By the induction hypothesis, the right-hand side vanish as $\delta\to 0$. Therefore, the desired limit in (B.4) holds in this case as well.
Now we complete the induction and establish (B.4).
Similar to [18, Lemma 6.3], we have
Lemma B.5.
Let $g\in C^{\infty}\left(\mathbb{C}^{m+n}\right)$ be independent of $z_{l}$, and ${\mathbf{i}},{\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m},{\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$. Then for all $r,s\in\mathbb{Z}_{\geq 0}$ such that $r+s<i_{l}$, and each $\delta>0$
  1. (1)
    $\int_{B_{\delta}^{\mathbf{j}}}z_{l}^{r}\bar{z}_{l}^{s}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=0;$
  2. (2)
    $\int_{S_{\delta}^{\mathbf{j}}}z_{l}^{r}\bar{z}_{l}^{s}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0$.
Here $d{\mathbf{z}}:=dz_{1}\wedge\cdots\wedge dz_{m+n}$ and $d\bar{\mathbf{z}}:=\overline{d{\mathbf{z}}}.$
Proof.
Let $B_{\delta,\hat{l}}^{{\mathbf{j}}}:=\Big{\{}{\mathbf{z}}_{\hat{l}}\in\mathbb{C}^{m+n-1}:\sup\{|z_{a}|:a\neq l\}<1,\delta<|{\mathbf{z}}_{\hat{l}}^{{\mathbf{j}}_{\hat{l}}}|<1\Big{\}}.$
For each ${\mathbf{z}}_{\hat{l}}\in B_{\delta,\hat{l}}^{\mathbf{j}}$, define
$B({\mathbf{z}}_{\hat{l}},\delta)=\left\{z_{l}\in\mathbb{C}:\delta/|{\mathbf{z}}_{\hat{l}}^{{{\mathbf{j}}}_{\hat{l}}}|<|z_{l}^{j_{l}}|<1\right\},$
The partial integrals
$I({\mathbf{z}}_{\hat{l}},\delta)$ $=\int_{B({\mathbf{z}}_{\hat{l}},\delta)}z_{l}^{r-i_{l}}\bar{z}_{l}^{s}\big{(}\ln(|z_{l}|^{2})\big{)}^{k_{l}}dz_{l}\wedge d\bar{z}_{l}$
$=\int_{\left(\delta/|{\mathbf{z}}_{\hat{l}}^{{\mathbf{j}}_{\hat{l}}}|\right)^{\frac{1}{j_{l}}}}^{1}\rho^{r-i_{l}+s+1}\big{(}\ln(\rho^{2})\big{)}^{k_{l}}d\rho\int_{0}^{2\pi}e^{i\theta\left(r-i_{l}-s\right)}d\theta$
are zero for all ${\mathbf{z}}_{\hat{l}}\in B_{\delta,\hat{l}}^{\mathbf{j}}$, because $r-s\leq r+s<i_{l}$. Then the integral in (1) is zero, since it is equal to
$\int_{B_{\delta,\hat{l}}^{{\mathbf{j}}}}I({\mathbf{z}}_{\hat{l}},\delta)\cdot{\mathbf{z}}_{\hat{l}}^{-{\mathbf{i}}_{\hat{l}}}\ln({\mathbf{z}}_{\hat{l}},{\mathbf{k}}_{\hat{l}})gd{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}$.
To compute (2), one integrates over $B_{\delta,\hat{l}}^{\mathbf{j}}$ the differential form $^{\mathbf{z}}_{\hat{l}}^{-{\mathbf{i}}_{\hat{l}}}\ln({\mathbf{z}}_{\hat{l}},{\mathbf{k}}_{\hat{l}})\cdot gd{\mathbf{z}}_{\hat{l}}$ $\wedge d\bar{{\mathbf{z}}}_{\hat{l}}$ times the integrals
$\int_{\left\{z_{l}:|z_{l}^{j_{l}}|=\delta/|{\mathbf{z}}_{\hat{l}}^{\mathbf{j}}|\right\}}z_{l}^{r-i_{l}}\bar{z}_{l}^{s}\big{(}\ln(|z_{l}|^{2})\big{)}^{k_{l}}d\bar{z}_{l},\quad{\mathbf{z}}_{\hat{l}}\in B_{\delta,\hat{l}}^{\mathbf{j}},$
which are all zero if $s\geq 0$, since in such case $r-i_{l}-s-1\leq r-i_{l}+s-1<0$. We deduce that the integral in (2) is zero. ∎
Similar to [18, Lemma 6.2],
Lemma B.6.
Let $k\in C^{\infty}(\mathbb{C}^{m+n})$, and ${\mathbf{i}}\in\mathbb{Z}_{\geq 0}^{m}$. There exists a decomposition
$k({\mathbf{z}},\bar{{\mathbf{z}}})=\sum_{r+s<i_{l}}z_{l}^{r}\bar{z}_{l}^{s}g_{r,s}^{l}({\mathbf{z}}_{\hat{l}},\bar{{\mathbf{z}}}_{\hat{l}})+K^{l}({\mathbf{z}},\bar{{\mathbf{z}}})$
$K^{l}({\mathbf{z}},\bar{{\mathbf{z}}})=\sum_{{\mathbf{k}}+{\mathbf{j}}={\mathbf{i}}}{\mathbf{z}}^{\mathbf{j}}\cdot\bar{{\mathbf{z}}}^{\mathbf{k}}\cdot K^{l}_{{\mathbf{j}},{\mathbf{k}}}({\mathbf{z}},\bar{{\mathbf{z}}})$
such that $g_{r,s}^{l}$ and $K_{{\mathbf{j}},{\mathbf{k}}}^{l}$ are smooth functions.
Proof.
The lemma is nearly identical to [18, Lemma 6.2], except that it does not involve summation over $l=1,\dots,m$. Consequently, the expression $g_{r,s}^{l}$ differs slightly from that in [18, Lemma 6.2].
The proof also follows similarly, based on a Taylor expansion.
Proceeding as in [18, Proposition 6.5], using B.4-B.6, we have:
Proposition B.7.
For any ${\mathbf{i}},{\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m},{\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$ and $g\in C^{\infty}\left(\mathbb{C}^{m+n}\right)$, the limit
$\lim_{\delta\rightarrow 0}\int_{B_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$ (B.5)
exists. Moreover,
$\lim_{\delta\rightarrow 0}\int_{S_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0.$ (B.6)
Proposition B.8.
Let $f$ be a never vanishing smooth function on $\bar{B}\subset\mathbb{C}^{m+n}$, and let ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m},{\mathbf{i}}\in\mathbb{Z}_{\geq 0}^{m}$. Denote
$S_{\delta}^{\mathbf{j}}(f)=\big{\{}{\mathbf{z}}\in B:|f{\mathbf{z}}^{\mathbf{j}}|=\delta\big{\}}$
$B_{\delta}^{\mathbf{j}}(f)=\big{\{}{\mathbf{z}}\in B:|f{\mathbf{z}}^{\mathbf{j}}|>\delta\big{\}}$
Consider $B_{\delta}^{\mathbf{j}}(f)$ with its canonical complex orientation and $S_{\delta}^{\mathbf{j}}(f)$ with the orientation such that $\partial B_{\delta}^{\mathbf{j}}(f)=-S_{\delta}^{\mathbf{j}}(f)$.
There exists an open neighborhood $\tilde{B}\subset B$ of the origin $0\in\mathbb{C}^{m+n}$ such that, if $\tilde{S}_{\delta}^{\mathbf{j}}:=S_{\delta}^{\mathbf{j}}(f)\cap\tilde{B}$ and $\tilde{B}_{\delta}^{\mathbf{j}}:=B_{\delta}^{\mathbf{j}}(f)\cap\tilde{B}$, then
$\lim_{\delta\rightarrow 0}\int_{\tilde{S}_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}=0$ (B.7)
for any $g\in C^{\infty}(\bar{B})$, and
$\lim_{\delta\rightarrow 0}\int_{\tilde{B}_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\lim_{\delta\rightarrow 0}\int_{B_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$ (B.8)
[…]
Proof.
We modify the argument in [18, Proposition 6.6], assuming that $|f|>1$ on $\bar{B}$.
Let $h$ be a smooth function such that $h^{j_{1}}=f$ on $\bar{B}$. Define a map $\mathbb{C}^{m+n}\to\mathbb{C}^{m+n}$ by
$\zeta_{1}=z_{1}h({\mathbf{z}},\bar{\mathbf{z}}),\quad\zeta_{l}=z_{l}\quad(l=2,\dotsc,n).$
Set
$\tilde{B}=\left\{{\mathbf{z}}\in B:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|<1\right\}.$
We may assume this map induces an isomorphism $\lambda:\tilde{B}\to B$.
Moreover, we have
$\lambda(\tilde{B}_{\delta}^{\mathbf{j}})=B_{\delta}^{\mathbf{j}},\quad\lambda(\tilde{S}_{\delta}^{\mathbf{j}})=S_{\delta}^{\mathbf{j}}.$
The inverse map $\mu:B\to\tilde{B}$ is given by
$z_{1}=\zeta_{1}\tilde{h}(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}}),\quad z_{j}=\zeta_{j}\quad(j=2,\dotsc,n),$
where $\tilde{h}$ is a smooth function on $B$ with no zeros.
Now take a function $k\in C^{\infty}(\bar{B})$. Changing coordinates via ${\mathbf{z}}=\mu(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}})$, we obtain:
$\ \ \ \ \int_{\tilde{S}_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})k({\mathbf{z}},\bar{{\mathbf{z}}})\,d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$ (B.9)
$=\int_{S_{\delta}^{\mathbf{j}}}{\boldsymbol{\zeta}}^{-{\mathbf{i}}}\sum_{l=0}^{i_{1}}\binom{i_{1}}{l}\left(\ln|\tilde{h}|^{2}\right)^{l}\left(\ln|\zeta_{1}|^{2}\right)^{i_{1}-l}\ln({\boldsymbol{\zeta}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})\tilde{k}({\boldsymbol{\zeta}},\bar{{\boldsymbol{\zeta}}})\,d\bar{\zeta}_{1}\wedge d{\boldsymbol{\zeta}}_{\hat{1}}\wedge d\bar{{\boldsymbol{\zeta}}}_{\hat{1}}$,
where
$\tilde{k}({\boldsymbol{\zeta}},\bar{{\boldsymbol{\zeta}}})=\tilde{h}^{-i_{1}}k(\mu,\bar{\mu})\cdot\Big{(}\bar{\tilde{h}}+\bar{\zeta}_{1}\frac{\partial\bar{\tilde{h}}}{\partial\bar{\zeta_{1}}}\Big{)}.$
It follows from B.7 and (B.9) that (B.7) holds.
To prove (B.8), we proceed by induction on $k_{1}$. Consider the case $k_{1}=0$. Thus,
[…]
$\frac{\partial b^{*}}{\partial z_{1}}=b.$
Indeed, set
$b^{*}(z):=\frac{1}{2\pi\sqrt{-1}}\int_{\mathbb{C}}\frac{b(\xi_{1}+z_{1},\bar{\xi}_{1}+\bar{z}_{1},{\mathbf{z}}_{\hat{1}},\bar{{\mathbf{z}}}_{\hat{1}})}{\bar{\xi}_{1}}\,d\xi_{1}\wedge d\bar{\xi}_{1},$ (B.10)
one can easily verify that $b^{*}$ satisfies the desired equation.
Set
$\gamma={\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b^{*}\,d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$. (B.11)
Then
$d\gamma={\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}.$ (B.12)
By Stokes’ theorem,
$\int_{\tilde{B}_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\int_{\partial\tilde{B}_{\delta}^{\mathbf{j}}}\gamma=\int_{{}^{1}\tilde{B}_{\delta}^{\mathbf{j}}}\gamma-\int_{\tilde{S}_{\delta}^{\mathbf{j}}}\gamma,$ (B.13)
where
$^{1}\tilde{B}_{\delta}^{\mathbf{j}}=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|=1,\sup_{a\neq 1}|z_{a}|<1,|{\mathbf{z}}_{\hat{1}}^{{\mathbf{j}}_{\hat{1}}}|>\delta\Big{\}}.$
In fact, to show the last equality of (B.13) , we use the decomposition
$\partial\tilde{B}_{\delta}^{\mathbf{j}}={}^{1}\tilde{B}_{\delta}^{\mathbf{j}}-\tilde{S}_{\delta}^{\mathbf{j}}+\sum_{l=2}^{n}{}^{l}\tilde{B}_{\delta}^{\mathbf{j}},$
with
$^{l}\tilde{B}_{\delta}^{\mathbf{j}}=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{1}h({\mathbf{z}},\bar{{\mathbf{z}}})|<1,|z_{l}|=1,\sup_{a\neq 1,l}|z_{a}|<1,|f{\mathbf{z}}^{\mathbf{j}}|>\delta\Big{\}},l\neq 1.$
It is clear that the integrals of $\gamma$ over the $^{l}\tilde{B}_{\delta}^{\mathbf{j}},l\neq 1$ vanish.
Consider
$U_{\delta}=\left\{{\mathbf{z}}\in B:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|>1,|{\mathbf{z}}_{\hat{1}}^{{\mathbf{j}}_{\hat{1}}}|>\delta\right\},$
so that
$\partial U_{\delta}=-{}^{1}\tilde{B}_{\delta}^{\mathbf{j}}+U_{\delta}^{\prime}+\sum_{l=1}^{n}{}^{l}U_{\delta},$
where
${}^{l}U_{\delta}$ $=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{l}|=1,\sup_{a\neq l}|z_{a}|<1,|{\mathbf{z}}_{\hat{l}}^{{\mathbf{j}}_{\hat{l}}}|>\delta,|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|>1\Big{\}},$
$U_{\delta}^{\prime}$ $=\left\{{\mathbf{z}}\in B:|{\mathbf{z}}_{\hat{1}}^{{\mathbf{j}}_{\hat{1}}}|=\delta,|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|>1\right\}.$
Again, the integrals of $\gamma$ over $U_{\delta}^{\prime}$ and $^{l}U_{\delta}$ for $l\neq 1$ vanish. Note that
$^{1}U_{\delta}={}^{1}B_{\delta}^{\mathbf{j}}:=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{1}|=1,\sup_{a\neq 1}|z_{a}|<1,|{\mathbf{z}}_{\hat{1}}^{{\mathbf{j}}_{\hat{1}}}|>\delta\Big{\}},$
since $|h|>1$ on $\bar{B}$, and the integral of $b$ over $U_{\delta}$ vanishes because $\mathrm{supp}(b)\subset\tilde{B}\subset\bar{B}\setminus U_{\delta}$. Hence, by Stokes’ theorem:
$0=\int_{\partial U_{\delta}}\gamma=\int_{{}^{1}\tilde{B}_{\delta}^{\mathbf{j}}}\gamma-\int_{{}^{1}B_{\delta}^{\mathbf{j}}}\gamma.$ (B.14)
Analogous to (B.13), we have
$\int_{B_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-\mathbf{i}}\ln({\mathbf{z}},\mathbf{k})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\int_{{}^{1}B_{\delta}^{\mathbf{j}}}\gamma-\int_{S_{\delta}^{\mathbf{j}}}\gamma.$ (B.15)
By (B.13), (B.14) and (B.15), we obtain:
$\left(\int_{B_{\delta}^{\mathbf{j}}}-\int_{\tilde{B}_{\delta}^{\mathbf{j}}}\right){\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\left(\int_{\tilde{S}_{\delta}^{\mathbf{j}}}-\int_{S_{\delta}^{\mathbf{j}}}\right)\gamma.$
By (B.6) and (B.7),
$\lim_{\delta\to 0}\left(\int_{\tilde{S}_{\delta}^{\mathbf{j}}}-\int_{S_{\delta}^{\mathbf{j}}}\right)\gamma=0.$
Therefore, (B.8) holds when $k_{1}=0$.
Now we consider the case $k_{1}>0$. Let $K=\mathrm{supp}(b)\subset\tilde{B}$ be the support of $b$. Then there exists $\alpha\in(0,1)$, s.t.
$K\subset K^{\prime}:=\big{\{}{\mathbf{z}}\in B:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|\leq\alpha,\sup\{|z_{a}|:a\neq 1\}\leq\alpha\big{\}}.$
Let
$K^{\prime\prime}:=\big{\{}{\mathbf{z}}\in B:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|\leq\alpha\big{\}}.$
[…]
We still take $b^{*}\in C^{\infty}(\bar{B})$ to be a smooth function given by (B.10) and define the (n,n1)(n,n-1)-form
$\gamma:={\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\,\phi b^{*}\,d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$. (B.16)
Then, (B.12) becomes
$d\gamma=\Big{(}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b+{\mathbf{z}}^{-\tilde{{\mathbf{i}}}}k_{1}(\ln|z_{1}|^{2})^{k_{1}-1}\ln({\mathbf{z}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})\phi b^{*}+{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\frac{\partial\phi}{\partial{z}_{1}}b^{*}\Big{)}d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}},$ (B.17)
where $\tilde{{\mathbf{i}}}=(i_{1}+1,i_{2},\cdots.i_{m}).$ Using (B.17) and proceeding as before, we obtain
$\ \ \ \ \lim_{\delta\to 0}\left(\int_{B_{\delta}^{{\mathbf{j}}}}-\int_{\tilde{B}_{\delta}^{{\mathbf{j}}}}\right){\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$
$=-\lim_{\delta\to 0}\left(\int_{B_{\delta}^{{\mathbf{j}}}}-\int_{\tilde{B}_{\delta}^{{\mathbf{j}}}}\right){\mathbf{z}}^{-\tilde{{\mathbf{i}}}}k_{1}(\ln|z_{1}|^{2})^{k_{1}-1}\cdot\phi b^{*}\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$
$\quad-\lim_{\delta\to 0}\left(\int_{B_{\delta}^{{\mathbf{j}}}}-\int_{\tilde{B}_{\delta}^{{\mathbf{j}}}}\right){\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\cdot\frac{\partial\phi}{\partial{z}_{1}}b^{*}\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$
$=:I_{1}+I_{2}.$
[…]
Proposition B.9.
Let ${\mathbf{i}},{\mathbf{k}},{\mathbf{l}}\in\mathbb{Z}_{\geq 0}^{m}$ and ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$. Then both limits below exist and are equal:
$\lim_{\delta\rightarrow 0}\int_{B_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\lim_{\delta\rightarrow 0}\int_{B_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$ (B.18)
for all $b\in C^{\infty}\left(\mathbb{C}^{m+n}\right)$.
Proof.
We have verified the existence of both limits. It now suffices to show that they are equal.
We proceed by induction on $m$ and $k_{1}$. When $m=1$, one observes that for each $\delta>0$, there exists $\delta^{\prime}>0$ such that $B_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}=B_{\delta^{\prime}}^{{\mathbf{j}}}$. In this case, it is clear that the two limits coincide.
Now assume $m>1$ and $k_{1}=0$.
Let $\gamma$ be the form as in (B.11). As in (B.13), by Stokes’ theorem,
$\int_{B_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\left(\int_{{}^{1}B^{{\mathbf{j}}+{\mathbf{l}}}_{\delta}}-\int_{S_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}}\right)\gamma.$ (B.19)
Moreover,
$\int_{{}^{1}B_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}}\gamma=\int_{B_{\delta,\hat{1}}^{{\mathbf{j}}+{\mathbf{l}}}}{\mathbf{z}}_{\hat{1}}^{-{\mathbf{i}}_{\hat{1}}}\ln({\mathbf{z}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})\,d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}\int_{|z_{1}|=1}z_{1}^{-i_{1}}b^{*}\,dz_{1}.$ (B.20)
By the induction hypothesis, the limit of this integral as $\delta\to 0$ is equal to
$\lim_{\delta\to 0}\int_{B_{\delta,\hat{1}}^{\mathbf{j}}}{\mathbf{z}}_{\hat{1}}^{-{\mathbf{i}}_{\hat{1}}}\ln({\mathbf{z}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}\int_{|z_{1}|=1}z_{1}^{-i_{1}}b^{*}dz_{1}=\lim_{\delta\to 0}\int_{{}^{1}B_{\delta}^{\mathbf{j}}}\gamma.$ (B.21)
Combining (B.15), (B.19)–(B.21) with (B.6), we obtain (B.18).
When $k_{1}>0$, we instead take $\gamma$ to be the form constructed as in (B.16). Then, applying the induction hypothesis as in B.8, we again obtain (B.18).
Remark B.10.
For any $\xi=(\xi_{1},\ldots,\xi_{n})\in\mathbb{C}^{n}$, define
$B(\xi):=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:\sup_{1\leq l\leq m}|z_{l}|<1,\ \sup_{1\leq l\leq n}|z_{m+l}-\xi_{l}|<1\Big{\}}\subset\mathbb{C}^{m+n}.$ (B.22)
For each ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$, define
$S^{\mathbf{j}}_{\delta}(\xi):=\{{\mathbf{z}}\in B(\xi):|{\mathbf{z}}^{\mathbf{j}}|=\delta\},\quad B_{\delta}^{\mathbf{j}}(\xi):=\{{\mathbf{z}}\in B(\xi):|{\mathbf{z}}^{\mathbf{j}}|\in(\delta,1)\}.$
Then all lemmas and propositions in this section extend to the setting of $S^{\mathbf{j}}_{\delta}(\xi)$ and $B_{\delta}^{\mathbf{j}}(\xi)$.
Proof of B.2.
We may as well assume that $|f|>1$ on $K.$
Let $K$ be the compact support of $\alpha$. Then $K$ can be covered by finitely many open sets $\{\tilde{B}_{a}\}_{a=1}^{N}$ together with an open set $V$, where each $\tilde{B}_{a}$ is a “box” centered at some point $(0,\xi)\in\mathbb{C}^{m}\times\mathbb{C}^{n}$ associated with $f$, chosen as in B.8, and $V$ is an open set satisfying
$V\cap\left\{{\mathbf{z}}\in\mathbb{C}^{n+m}:|f{\mathbf{z}}^{{\mathbf{j}}}|<\delta_{0}\right\}=\emptyset$
for some $\delta_{0}>0$. Moreover, we may assume without loss of generality that $f\neq 0$ on the closure of $\left(\bigcup_{a=1}^{N}\tilde{B}_{a}\right)\cup V.$
Let $\{\phi_{a}\}_{a=1}^{N+1}$ be a partition of unity subordinate to the open cover $\left(\bigcup_{a=1}^{N}\tilde{B}_{a}\right)\cup V$, with $\phi_{N+1}$ supported inside $V$.
It then follows from (B.8) and B.10 that
$\lim_{\delta\to 0}\sum_{a=1}^{N}\int_{|f{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\phi_{a}\alpha=\lim_{\delta\to 0}\sum_{a=1}^{N}\int_{|{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\phi_{a}\alpha.$ (B.23)
Since $f\neq 0$ on $K$, there exists $\delta_{0}^{\prime}>0$ such that
$V\cap\left\{{\mathbf{z}}\in\mathbb{C}^{n+m}:|{\mathbf{z}}^{{\mathbf{j}}}|<\delta_{0}^{\prime}\right\}=\emptyset.$
Then for any $\delta<\min\{\delta_{0},\delta_{0}^{\prime}\}$, we have
$\int_{|f{\mathbf{z}}^{{\mathbf{j}}}|<\delta}\phi_{N+1}\alpha=\int_{|{\mathbf{z}}^{{\mathbf{j}}}|<\delta}\phi_{N+1}\alpha.$ (B.24)
By combining (B.23) and (B.24), we conclude that
$\lim_{\delta\to 0}\int_{|f{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\alpha=\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\alpha.$ (B.25)
Now let ${\mathbf{j}}^{\prime}=(j_{1}^{\prime},\ldots,j_{m}^{\prime})\in\mathbb{Z}_{>0}^{m}$, and define
$\tilde{{\mathbf{j}}}:=\left(\max\{j_{1},j_{1}^{\prime}\},\ldots,\max\{j_{m},j_{m}^{\prime}\}\right).$
Then, by applying a similar partition of unity argument and using B.9, we obtain
$\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\alpha=\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{\tilde{{\mathbf{j}}}}|>\delta}\alpha=\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{{\mathbf{j}}^{\prime}}|>\delta}\alpha.$ (B.26)
It follows from (B.25) and (B.26) that the limit in (B.1) is independent of the choices of $f$ and $\mathbf{j}$}.
Global case
Now, we generalize Cauchy principal values to complex manifolds.
Definition B.11.
Let $M$ be a complex manifold with a simple normal crossing divisor $D$. A coordinate chart compatible with $D$ is an open subset $U\subset M$ together with a biholomorphic map
$\varphi:U\rightarrow V\subset\mathbb{C}^{m+n},$
where $V$ is an open subset of $\mathbb{C}^{m+n}$, and $m,n$ are non-negative integers, such that $\varphi(D\cap U)=(\{0\}\times\mathbb{C}^{n})\cap V$. We say that a property holds locally, if it holds in any coordinate chart compatible with $D$.
Definition B.12.
[…]
Remark B.13.
For holomorphic vector bundles over complex manifolds with simple normal crossing divisors, we can also define sections with divisorial type singularities in a similar manner.
One nice property of divisorial type singularities is that they are stable under pull back:
Proposition B.14.
Let $M$ and $N$ be complex manifolds with simple normal crossing divisors $D$ and $D^{\prime}$ respectively. Assume $F:M\rightarrow N$ is a holomorphic map which satisfies $F^{-1}(D^{\prime})=D$ as sets, and $\alpha$ is a differential form on $N$ with divisorial type singularities along $D^{\prime}$. Then $F^{*}(\alpha)$ is a differential form on $M$ with divisorial type singularities along $D$. Moreover, if $\alpha$ is logarithm-free, $F^{*}(\alpha)$ is also logarithm-free.
Proof.
We only need to check that $F^{*}(\alpha)$ has divisorial type singularities (possible logarithm-free) locally. The local case can be derived from the following fact: for a open subset $U\subset\mathbb{C}^{m+n}$ and a holomorphic function $g:U\rightarrow\mathbb{C}$, which satisfies
$g^{-1}(0)=\{(z_{1},z_{2},...,z_{m+n})\in U:z_{i}=0\text{ for some }i\in\{1,2,...,m\}\},$
there exist ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$ and a nowhere vanishing holomorphic function $\tilde{g}$ on $U$, such that $g=\tilde{g}\,{\mathbf{z}}^{{\mathbf{j}}}$. ∎
Now, we can define the Cauchy principal value.
Definition B.15.
Let $M$ be a complex manifold with a simple normal crossing divisor $D$. Assume $\{U_{i}\}_{i\in I}$ is a coordinate cover compatible with $D$, and $\{\rho_{i}\}_{i\in I}$ is a partition of unity for $\{U_{i}\}_{i\in I}$. Given a compactly supported differential form with divisorial type singularities along $D$, which we denote by $\alpha$. The integral of $\alpha$ over $M$, in the sense of Cauchy principal value, is defined by
[…]
It is standard to prove the Cauchy principal value does not depend on the choice of $\{U_{i}\}_{i\in I}$ and $\{\rho_{i}\}_{i\in I}$.
Finally, we introduce another way to characterize the Cauchy principal value on complex manifolds, which will be used in the main content.
Definition B.16.
Let $M$ be a complex manifold with simple normal crossing divisor $D$. A smooth defining function of $D$ is a smooth function $h:M\rightarrow\mathbb{R}$ that satisfies the following conditions:
  1. (1)
    $h$ is a non-negative, and $h^{-1}(0)=D$.
  2. (2)
    For each point $p\in D$, there exists an open neighborhood $U\subset M$ of $p$, a smooth positive function $f:U\rightarrow\mathbb{R}$, and a holomorphic function $g:U\rightarrow\mathbb{C}$, such that $h^{-1}(0)=D\cap U$ and $h|_{U}=fg\bar{g}$.
The concept of smooth defining function is stable under pull back:
Proposition B.17.
Let $M$ and $N$ be complex manifolds with simple normal crossing divisors $D$ and $D^{\prime}$ respectively. Assume $F:M\rightarrow N$ is a holomorphic map which satisfies $F^{-1}(D^{\prime})=D$ as sets, and $h$ is a smooth defining function of $D^{\prime}$. Then $h\circ F$ is a smooth defining function of $D$.
Proof.
This can be verified directly. ∎
The following proposition provides another way to characterize the Cauchy principal value:
Proposition B.18.
Let $M$ be a complex manifold with a simple normal crossing divisor $D$. Let $h$ be a smooth defining function of $D$, and let $\alpha$ be a compactly supported differential form with divisorial-type singularities along $D$. Then we have
$\lim_{\delta\rightarrow 0}\int_{h>\delta}\alpha=\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{M}\alpha.$
Proof.
[…]
We notice that $\mathrm{supp}(\rho_{0}\alpha)$ is compact, so $h|_{\mathrm{supp}(\rho_{0}\alpha)}$ has a nonzero minimum. As a consequence,
$\lim_{\delta\rightarrow 0}\int_{h>\delta}\rho_{0}\alpha=\int_{M}\rho_{0}\alpha=\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{M}\rho_{0}\alpha.$ (B.27)
For $\rho_{p}\alpha$, we notice the following fact: for a open subset $U\subset\mathbb{C}^{m+n}$ and a holomorphic function $g:U\rightarrow\mathbb{C}$, which satisfies
$g^{-1}(0)=\big{\{}(z_{1},z_{2},\cdots,z_{m+n})\in U:z_{i}=0\text{ for some }i\in\{1,2,...,m\}\big{\}},$
there exist ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$ and a nowhere vanishing holomorphic function $\tilde{g}$ on $U$, such that $g=\tilde{g}\,{\mathbf{z}}^{{\mathbf{j}}}$. By using this fact and Theorem B.2, we have
$\lim_{\delta\rightarrow 0}\int_{h>\delta}\rho_{p}\alpha=\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{M}\rho_{p}\alpha$ (B.28)
for $p\in D$. Combining (B.27) and (B.28), we proved this proposition. ∎